### v1

# Evaluation of Language Models in the Medical Context Under Resource-Constrained Settings

Abstract
Since the emergence of the Transformer architecture, language model development has increased, driven by their promising potential. However, releasing these models into production requires properly understanding their behavior, particularly in sensitive domains such as medicine. Despite this need, the medical literature still lacks technical assessments of pre-trained language models, which are especially valuable in resource-constrained settings in terms of computational power or limited budget. To address this gap, we provide a comprehensive survey of language models in the medical domain. In addition, we selected a subset of these models for thorough evaluation, focusing on classification and text generation tasks. Our subset encompasses 53 models, ranging from 110 million to 13 billion parameters, spanning the three families of Transformer-based models and from diverse knowledge domains. This study employs a series of approaches for text classification together with zero-shot prompting instead of model training or fine-tuning, which closely resembles the limited resource setting in which many users of language models find themselves. Encouragingly, our findings reveal remarkable performance across various tasks and datasets, underscoring the latent potential of certain models to contain medical knowledge, even without domain specialization. Consequently, our study advocates for further exploration of model applications in medical contexts, particularly in resource-constrained settings. The code is available on https://github.com/anpoc/Language-models-in-medicine.
1 Introduction
Natural Language Processing (NLP) holds immense promise in the medical domain. Recently, the medical community has shown substantial interest in leveraging state-of-the-art language models to address various medical challenges [1, 2]. In particular, generative large language models (LLMs) have showcased emergent abilities beyond their original training objectives, such as text summarization and question-answering [3]. These newfound abilities have enabled LLMs to perform tasks of significant clinical importance, including passing medical examinations, summarizing clinical and radiological reports, as well as medical dialogues, extracting drug names from medical notes, responding to patient inquiries, and writing medical histories and physical assessments [2, 4].
The versatility of language models can be attributed to a convergence of factors [2, 4, 5]. The first factor is their ability to learn valuable patterns within large amounts of unlabeled data via self-supervision. The second factor revolves around the Transformer architecture [6] and its suitability for efficient parallel processing on modern computing hardware. Lastly, the third factor encompasses the crucial process of fine-tuning language models to align their responses with human expectations through instruction tuning.
Integration of language models in medical settings is becoming a reality as partnerships between developers and healthcare systems continue to grow [7]. The potential benefits are significant, as they can derive broadly applicable representations from extensive medical corpora at scale. Recent research indicates that LLMs with more than 100 billion parameters can encapsulate clinical knowledge [8]. Nevertheless, it is essential to recognize that our understanding of the behavior of LLMs and language models in general still needs to be completed [4]. Moreover, implementing these models also carries risks, such as the generation of inaccurate results, a phenomenon known as hallucinations, and the potential amplification of existing biases [1, 4].
Hence, the deployment of language models in sensitive fields, such as healthcare, should be approached with the utmost care [5]. To further understand and assess language models’ performance in clinical scenarios, we conducted comprehensive evaluations using models of varying sizes, ranging from 110M to 13B parameters, and encompassing different knowledge domains. We focus on text classification, using three different approaches and conditional text generation tasks. The approaches employed for text classification and the application of zero-shot prompting alongside allow performance improvement without the need for training or fine-tuning. For its part, the selection of the model size range considers resource constraints, in terms of limited budgets or computational power, encountered not only by many research groups but also by other actors and institutions, such as hospitals, where handling models with more than 110 billion parameters is not feasible.
Moreover, the use of small language models aligns with the crucial imperative of environmental sustainability, as larger models often require significant energy and computational resources. It also opens up possibilities for organizations to develop applications that can run directly on commodity hardware and small devices rather than relying on cloud-based services. In fact, the prospect of deploying these language models on smartphones and other mobile devices operating at the edge without cloud dependence is promising [9]. Therefore, the use of language models in resource-constrained settings not only addresses practical challenges but also has great potential in edge and local computing.
The paper is structured as follows. Section 1 introduces the study and provides an overview of its objectives. Section 2 elaborates on the preliminaries, establishing foundational concepts relevant to the investigation. Related work in the existing literature is explored in Section 3. Section 4 details the methodology adopted, and Section 5 provides the most significant results and analysis. Finally, Section 6 contains the conclusions drawn from the study, summarizing key insights and implications for future research directions.
2 Preliminaries
The evolution of natural language processing can be condensed into four major groups of models: (1) statistical models, (2) neural language models, (3) pre-trained language models, and (4) large language models [10]. Each of these groups represents a paradigm shift in natural language modeling and has contributed significantly to the conception of language models as we know them today.
The first transition, from statistical to neural language models, marked a shift from predicting words based on minimal local context to probabilistically assessing word sequences using neural networks. Additionally, this transition introduced the concept of representing words as low-dimensional continuous embeddings based on their contextual usage (distributional semantics). The second transition, from neural language models to pre-trained language models, represented a move from task-specific model development towards pre-training and fine-tuning methodologies. Lastly, the third transition to large language models constituted a transformation from discriminative AI to generative AI, from model-centric to data-centric approaches, and from fine-tuning to prompt-tuning [10, 11, 12]. These advances have paved the way for more sophisticated language models with broader applications and improved capabilities.
2.1 From task-specific model development towards pre-training and fine-tuning: pre-trained language models
The emergence of pre-trained language models gave relevance to pre-training and fine-tuning methodologies, which was a paradigm shift, as mentioned above. This shift has driven research toward designing more efficient architectures and refinement of pre-training strategies, demonstrating significant advantages in enriching language understanding and improving model performance in various tasks [11, 10]. One of the first attempts at pre-trained language models is ELMo [13]. ELMo was developed to capture context-aware word representations by pre-training a bidirectional Long Short-Term Memory (biLSTM) network and fine-tuning it for subsequent downstream tasks. However, the introduction of the Transformer architecture represented a qualitative leap in NLP, offering highly parallelizable structures and self-attention mechanisms.
The Transformer [6] follows the autoencoder archetype, giving rise to three families of models: (1) BERT-family or encoder-only models, (2) GPT-family or decoder-only models, and (3) text-to-text or encoder-decoder models. Their graphical representations are shown in Fig. 1.
2.1.1 Encoder-only models
Encoder-only models, exemplified by BERT [14], are based on masked language modeling (MLM), where parts of the input are masked to encourage the model to reconstruct the original sequence, leveraging contextual information bidirectionally. These models can be stated as . In particular, their contextual embeddings have been proven highly effective as general-purpose semantic features, significantly boosting performance in discriminative NLP tasks.
2.1.2 Decoder-only models
Decoder-only models focus on autoregressive language modeling, that is, predicting the next token in a sequence based on previous tokens. These models produce contextual embeddings and a distribution over the next tokens , which can be stated as . However, the contextual embeddings they generate depend solely on the left context. Currently, most research efforts are directed toward decoder-only models due to their exceptional performance in conditional generation tasks and their demonstrated emergent capabilities.
2.1.3 Encoder-decoder models
Text-to-text models, or encoder-decoder models, are trained to learn the correspondence between a pair of texts and can be stated as . These models combine bidirectional contextual embeddings with the capability to generate output sequences, making them versatile in various text-to-text tasks without requiring additional heads for fine-tuning. Moreover, by having a wide spectrum of language tasks that can be translated into text-to-text representation, these models can potentially be used for a wide range of applications.
2.2 From discriminative AI to generative AI: large language models
More recently, it has been observed that increasing the size of language models, whether in terms of model size, dataset size, or computational resources, often results in improved model capabilities in various tasks [15, 16, 17, 18, 19, 20, 21], including those that require specialized scientific knowledge and reasoning [22]. Known as the scaling law [23], this phenomenon posits a power-law relationship between model performance, measured by the loss, and the factors mentioned above.
Several studies have delved into this relationship by progressively training larger language models, up to hundreds of billion parameters, such as GPT-3 [17], PaLM [19], Galactica [24], LLaMA models [25, 26], Claude [27], Gemini 1.5 [28], and Mistral [29]. Although scaling primarily involves increasing model size while retaining similar architectures and pre-training tasks as smaller models [10], large language models (LLMs) have exhibited unexpected emergent capabilities [30, 3], enabling them to address a wide range of complex tasks. GPT-3 and ChatGPT serve as precursors to these large language models [11, 10].
A particularly intriguing aspect of LLMs is their few-shot and zero-shot capabilities [31], enabling them to adapt to various tasks without requiring gradient-based parameter updates [8, 12, 32, 17, 21, 33, 34, 35]. Through these and other prompting techniques, emerging capabilities have been observed, enabling LLMs to exhibit apparent reasoning abilities [8]. Nonetheless, prompting techniques are not exclusive to LLMs but are also applicable, for example, to pre-trained language models. Similarly, numerous studies highlight the competence of LLMs as implicit knowledge bases [8, 22, 24, 36]. As for prompt engineering, its role is highly significant due to its benefits, which have proven to be more effective in large or instruction-tuned models. [21, 32, 8]. In the medical domain, for example, Chain of Thought (CoT) has been used to address explainability [37] and in-context learning to mitigate the need for costly medical annotations [11].
Among LLMs, GPT-4 stands out for its exceptional performance, often matching or surpassing human performance in various tasks [38, 39, 40], including in scientific domains such as biology, chemistry, and medicine [41]. In the general domain, extensive evaluations of this model have been conducted [22, 42, 43, 44], even exploring the path toward Artificial General Intelligence (AGI) [39]. In the medical domain, recent studies have begun to investigate its utility in various applications, ranging from medical chatbots [45] to medical competency exams [40] and its applications in radiology [46, 38, 47, 48, 49, 50, 51], among others [52, 53, 54, 55].
Nevertheless, there are other relevant models in the medical field, and unlike GPT-4, most of them are domain-specific. Generally, these models are derived from general-purpose LLMs, with exceptions like GatorTron [56], Galactica, and GatorTronGPT [57]. The most common training technique is instruction fine-tuning, as seen in Galactica, MedAlpaca [58], PMC-LLaMA [59], Med-PaLM 2 [60], GatorTronGPT, and ClinicalGPT [61]. However, although less frequently, reinforcement learning from human feedback (RLHF) and reinforcement learning from AI feedback (RLAIF) techniques are also adopted, HuatuoGPT [62] being an example. Recent research studies also indicate a multimodal trend that supports various types of healthcare data, including electronic health records (EHR), medical images, and medical sequence signals. Examples of these developments include LLaVAMed [63], MedAGI [64], OphGLM [65], Visual Med-Alpaca [66], and MedFlamingo [67], CheXzero [68].
Despite their advantages, LLMs also have limitations. One of these are computational challenges due to their high computational resource requirements [8, 10]. Conducting repetitive and ablation studies to investigate the impact of training strategies for LLMs is costly and time-consuming. Therefore, in certain cases, LLMs may not represent a cost-effective solution. Since smaller language models have also shown the effectiveness of using curated scientific and biomedical corpora for both discriminative and generative language modeling [38], they represent an alternative to LLMs. Such models include BioBERT [69], SciBERT [70], PubMedBERT [71], BioMegatron [72], ScholarBERT [73], BioGPT [74], and ClinicalBERT [75]. However, the current literature still lacks comparative studies on these smaller language models.
3 Related Work
Comparative studies investigating language models are crucial in advancing our understanding of them, shedding light on their functionalities, and pinpointing their constraints. Despite previous research, a notable gap persists in the literature. This gap is particularly significant in fields that require heightened sensitivity, such as medicine, where a thorough understanding of models is imperative. Although existing research in this domain focuses primarily on specific tasks, datasets, or models [5, 32, 12, 38, 8], there are exceptions. Some of these studies are detailed below.
In [11], the authors conduct a descriptive study delineating the capabilities and limitations of language models within the healthcare domain, encompassing pre-trained and large language models. They explain the development process, provide insight into training data, methodologies, and optimization strategies, and investigate concerns about integrating LLMs in healthcare. These concerns focus on fairness, accountability, transparency, and ethics. Likewise, in the context of integration concerns, the faithfulness problem within existing research on artificial intelligence in healthcare and medicine is explored in [76]. This analysis involves identifying the causes of nonfaithful results, possible evaluation metrics for faithfulness, and potential mitigation strategies for this issue.
Regarding studies closely aligned with the present work, notable contributions include those by Soni et al. [77], Lehman et al. [78], and Zhou et al. [79]. In [77], the authors assess the cost-effectiveness of pre-training and fine-tuning in Transformer language models, specifically in BERT, BioBERT [69], Clinical BERT [75], and XLNet [80]. These models underwent pre-training and fine-tuning on general, biomedical, and clinical-domain data combinations for medical question-answering (QA) tasks. Their findings revealed that BERT-based models exhibit superior performance when fine-tuned on multiple datasets, incorporating both general and clinical domain data. This observation suggests a gap in well-generalizable medical QA datasets. Furthermore, their results underscored the positive impact of initial fine-tuning on general-domain datasets, such as SQuAD, before fine-tuning on clinical datasets, highlighting the quality of the large, manually constructed SQuAD dataset. Overall, it was observed that initial fine-tuning improves performance in most cases. It is important to note that this study focused solely on question-answering and did not evaluate prompting techniques.
On its side, in [78], the authors’ research delves into whether LLMs, primarily trained on general web text, are suitable for highly specialized, safety-critical domains such as medicine or if domain-specialized models offer better performance. The authors evaluated 12 language models, ranging from 220 million to 175 billion parameters, measuring their performance in three clinical tasks. As part of their experiments, T5 models were trained from scratch using MIMIC-III and MIMIC-IV clinical notes to investigate the efficiency of clinical tokens. Their findings suggest that relatively small specialized clinical models significantly outperform all in-context learning approaches, even when fine-tuned on limited annotated data. However, this study does not assess the ability of models to handle long text and considered neither decoder-only models nor instruction-tuned ones.
Lastly, [79] provides a comprehensive overview of the development and deployment of LLMs in medicine, including the challenges and opportunities that they face. The main highlight of this study is that it is not only a discursive but also a practical assessment. The authors begin with an introduction to the principles of existing medical LLMs, including their basic model structures, the number of parameters, and the sources and scales of data used for model development. A comparative evaluation of the performance of different LLMs in various medical tasks is then provided, comparing them with state-of-the-art lightweight models. This review is framed by the research questions: 1) What are the practices for developing medical LLMs? 2) How to measure the medical task performance of LLMs in a medical setting? 3) How have medical LLMs been used in actual practice? 4) What are the challenges in using medical LLMs; and 5) How to develop and deploy medical LLMs more effectively?
4 Methodology
The experiments involve two target tasks: text classification and conditional text generation. Specifically, text classification is addressed through three different approaches: (i) contextual embedding similarity, (ii) natural language inference (NLI), and (iii) multiple-choice question-answering (MCQA). These are explained in Section 4.1.2. As detailed below, each task is evaluated using diverse datasets, models, and performance metrics.
The evaluation process comprises 53 models, which exhibit a size spectrum ranging from 110 million to 13 billion parameters. These models belong to three knowledge domains: general, biomedical or scientific, and clinical. Their characterization can be found in Table 1. OpenLLaMA 13B, Flan-T5-XXL, T5-V1.1-11B, and T0++ are run using float16 precision.
In order to conduct a comprehensive analysis, we recognize the importance of evaluating the chosen models across a broad spectrum of medical specialties, ultimately aiming at a general assessment of medical knowledge. To achieve this, we used the transcriptions dataset. Nevertheless, also recognizing the potential of artificial intelligence in particular medical fields, we narrowed our focus in addition to radiology, one of the most promising fields for AI integration. Consequently, we opt for the MIMIC-CXR dataset along with its labeled counterpart, MS-CXR.
| Small (S) | Medium (M) | Large (L) | XL | XXL | ||||||||||||
| ID | Model | Size | ID | Model | Size | ID | Model | Size | ID | Model | Size | ID | Model | Size | ||
| Encoder-only | General | m00 | BERT 0 [14] | 110 M | m01 | BERT 0 [14] | 340 M | - | - | - | - | - | - | - | - | - |
| Biomedical | m02 | BiomedBERT | 110 M | m04 | BiomedBERT-large | 340 M | - | - | - | - | - | - | - | - | - | |
| (abstracts + full text) 0 [71] | (abstracts only) 0 [71] | |||||||||||||||
| m03 | BiomedBERT | 110 M | m12 | RoBERTa-MNLI ‡1 [81] | 355 M | |||||||||||
| (abstracts only) 0 [71] | ||||||||||||||||
| m05 | SciBERT 0 [70] | 110 M | ||||||||||||||
| m06 | SapBERT 0 [82] | 110 M | ||||||||||||||
| m07 | BioLORD-STAMB2-v1 0 [83] | 110 M | ||||||||||||||
| m08 | BioLORD-STAMB2-v1-STS2 0 [83] | 110 M | ||||||||||||||
| m09 | BioLORD-PMB 0 [83] | 110 M | ||||||||||||||
| m11 | NLI-DeBERTa ‡1 [84] | 100 M | ||||||||||||||
| Clinical | m10 | Bio+Clinical BERT 0 [75] | 110 M | - | - | - | - | - | - | - | - | - | - | - | - | |
Encoder-
decoder |
General | m14 | T5-V1.1-Base 2 [85, 86] | 220 M | m13 | BART Large-MNLI 1 [87] | 407 M | m15 | T5-V1.1-Large 2 [85, 86] | 770 M | m16 | T5-V1.1-3B 2 [85, 86] | 3.0 B | m17 | T5-V1.1-11B 2 [85, 86] | 11.0 B |
| m18 | Flan-T5-Base †2 [15] | 220 M | m19 | Flan-T5-Large †2 [15] | 770 M | m20 | Flan-T5-XL †2 [15] | 3.0 B | m21 | Flan-T5-XLL †2 [15] | 11.0 B | |||||
| m22 | T0 3B †2 [88] | 3.0 B | m23 | T0++ †2 [88] | 11.0 B | |||||||||||
| Biomedical | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | |
| Clinical | m24 | ClinicalT5-base 2 [89] | 220 M | - | - | - | m25 | ClinicalT5-large 2 [89] | 700 M | - | - | - | - | - | - | |
| Decoder-only | General | - | - | - | m26 | GPT-2 Medium3 [90] | 355 M | m27 | GPT-2 Large 3 [90] | 774 M | m28 | GPT-2 XL 3 [90] | 1.5 B | m29 | Palmyra Base 5B 23 [91] | 5.0 B |
| m41 | OpenLLaMA 3B 3 [92] | 3.0 B | m30 | Camel 5B†2 [93] | 5.0 B | |||||||||||
| m42 | OpenLLaMA 3Bv2 3 [92] | 3.0 B | m31 | GPT-J 6B 23 [94] | 6.0 B | |||||||||||
| m32 | Instruct GPT-J †2 [95] | 6.0 B | ||||||||||||||
| m33 | Falcon-7B 23 [96] | 7.0 B | ||||||||||||||
| m34 | Falcon-7B-Instruct †2 [96] | 7.0 B | ||||||||||||||
| m35 | MPT-7B 23 [97] | 7.0 B | ||||||||||||||
| m36 | MPT-7B-Instruct †2 [97] | 7.0 B | ||||||||||||||
| m37 | LLaMA-7B 23 [25] | 7.0 B | ||||||||||||||
| m38 | LLaMA 2-7B 23 [26] | 7.0 B | ||||||||||||||
| m39 | Alpaca 7B †2 [98] | 7.0 B | ||||||||||||||
| m40 | LLaMA 2-CHAT-7B †2 [26] | 7.0 B | ||||||||||||||
| m43 | OpenLLaMA 7B 3 [92] | 7.0 B | ||||||||||||||
| m44 | OpenLLaMA 7Bv2 3 [92] | 7.0 B | ||||||||||||||
| m45 | OpenLLaMA 13B 3 [92] | 13.0 B | ||||||||||||||
Biomedical
/ Scientific |
- | - | - | m48 | BioGPT 3 [74] | 347 M | m47 | GPT-2-PubMed Large 3 [99] | 774 M | m50 | Galactica 1.3B 3 [24] | 1.3 B | m51 | Galactica 6.7B 3 [24] | 6.7 B | |
| m46 | GPT-2-PubMed Medium 3 [99] | 355 M | m49 | BioGPT-Large 3 [74] | 1.5 B | |||||||||||
| Clinical | - | - | - | - | - | - | - | - | - | - | - | - | m52 | MedAlpaca 7b †2 [58] | 7.0 B | |
4.1 Text classification
For the text classification task, we evaluate two datasets using three approaches and their corresponding models. It is essential to mention that fine-tuning is not performed; however, we use zero-shot prompt-engineering. In the following, we provide a concise overview of the datasets, followed by an explanation of the approaches employed. We refer the reader to Appendix A for more details on the preprocessing steps and the data characteristics of the datasets.
4.1.1 Datasets
Transcriptions is a multi-label collection of electronic health records covering several medical specialties. The preprocessing procedure for this dataset involves removing null entries, modifying the report’s format, and selecting and renaming labels. After preprocessing, the resultant dataset consists of 2074 samples and 29 labels. Model performance is evaluated using the AUC score.
In particular, since the EHRs are used in their entirety, certain samples contain token vectors that exceed the maximum input length allowed for multiple models. There are even samples whose token vectors exceed 2048 tokens, the maximum input length supported among the models. Consequently, different strategies have been adopted for each approach to effectively address this challenge. The adopted strategies, which require splitting the sentence into parts, will produce different results than processing the entire sample at once due to the attention mechanisms. These approaches are detailed in Section 4.1.2.
MS-CXR is a multi-class dataset comprising sections of X-ray reports, each accompanied by annotations made by a radiologist [100, 101, 102]. The number of classes is eight, well distributed and represented in 718 unique text samples. The preprocessing procedure is more straightforward than for the other datasets and includes removing samples with missing information and duplicates. In contrast to transcriptions, there is no evidence of any sample exceeding the maximum input length allowed for any of the models.
Model performance is evaluated using a set of metrics in their macro-averaged version to ensure a comprehensive assessment. The metrics are accuracy, F1-score, precision, and recall.
4.1.2 Approaches
Contextual embedding similarity is grounded in the cosine similarity between the contextual embeddings of the text under consideration and its corresponding label. The models employed are exclusively encoder-only, comprising 11 models subjected to evaluation. The final contextual or sentence embedding is determined using three distinct pooling strategies: the CLS-token embedding, average pooling of the token-level embedding, and maximum pooling of the token-level embedding. In addition, prompt-engineering is implemented for the label. The specific details of the proposed prompts can be found in Appendix B.
In cases where the input tokens exceed the maximum input size defined by the model, the samples are split into smaller parts, processing each individually. The embedding of the sample varies depending on the pooling strategy implemented, as described below. In particular, the models used in this approach have a maximum input token size of 512 tokens.
CLS pooling: The contextual embeddings of the constituent parts of a sample are obtained by CLS pooling. The contextual embedding of the sample is calculated as the element-wise average among the contextual embeddings of the constituent parts.
Maximum pooling: The contextual embedding of each sample is determined as the element-wise maximum of the output token-level embeddings of the constituent parts.
Average pooling: The average operator considers the number of non-padding tokens of the parts into which a sample’s text is segmented. Thus, the contextual embedding of each sample is calculated as the element-wise weighted average among the token-level embeddings of its constitutive parts.
Natural language inference for text classification involves representing the text to be classified as a premise and the labels as hypotheses. Subsequently, a multi-classification task with three classes -entailment, neutral, and contradiction- is performed. Conceptually, this approach can be thought of as having classifiers, where corresponds to the number of labels or classes. The models employed are encoder-only (cross-encoders) or encoder-decoder architectures.
How the output is processed varies depending on whether the dataset is multi-label or multi-class. The entailment and contradiction outputs are transformed into binary probabilities in the multi-label scenario. This transformation effectively translates the problem into binary text classifiers, each classifier for one of the labels. In contrast, in the multi-class scenario, only the entailment logits are used to predict the class associated with a sample. When the length of the input tokens exceeds the maximum input size defined by the model, the scores of these samples are calculated by aggregating the scores obtained from their constituent parts. Aggregation measures that can be considered include maximum, average, or median.
Furthermore, like the previous approach, the influence of prompts is considered during the model evaluation. The prompts used in this context are identical to those employed in the contextual embedding similarity approach, and their details can be found in Appendix B.
Multiple-choice question-answering framework enables encoder-decoder and decoder-only architectures to perform text classification. A total of 27 models are assessed in this approach. Table 1, which provides details of the models used, shows the inclusion of pre-trained models and their instruction-tuned counterparts.
For this approach to work, a prompt is required. Therefore, the impact of the prompt is measured in the analysis. The prompts are formulated based on those most commonly used in instruction-tuning models for multiple-choice question-answering tasks. However, in these prompts, the question format is not restrained. For this reason, this study also considers the impact of the question itself. Additional information on the prompts used in this approach can be found in Appendix B.
It is essential to mention that the logit space has been constrained to align with the set of viable alternatives presented in a multiple-choice scenario. For this, the token identifiers associated with the feasible response options are determined and used to filter the logit space.
4.2 Conditional text generation
We evaluate one dataset for the conditional text generation task, using perplexity as the evaluation metric. Perplexity is a measure of uncertainty on the value of a sample from a discrete probability distribution. If we have a tokenized sequence , then
where is the log-likelihood of the -th token conditioned on the preceding tokens according to our model.
Below, a concise overview of the dataset employed is provided, followed by an explanation of the approach adopted. A more detailed characterization of MIMIC-CXR and its preprocessing procedure is presented in Appendix A.
4.2.1 Datasets
MIMIC-CXR is composed of X-ray reports [103, 104, 102]. As part of the preprocessing procedure, the relevant sections are extracted using the code provided by Johnson et al. [105, 106]. Following this extraction process and the subsequent removal of null and duplicated samples, the resulting dataset contains samples. Furthermore, none of the samples exceed the maximum input size allowed for the proposed models.
4.2.2 Approaches
For this task, 20 decoder-only models are evaluated, as detailed in Table 1. Among these models, we find Galactica, whose tokenizer lacks special tokens. As a result, the analysis is conducted in two parts. In the first part, the Beginning of Sequence (BOS) token is included during tokenization, with perplexity calculated from the first token of the texts. The BOS token is a special token usually employed by generative models to indicate the start of a text. In cases where a model’s tokenizer does not have the BOS token predefined, such as the Falcon-7B, it is defined as its first special token. The second part of the analysis proceeds without the BOS token, and perplexity is determined from the second token of the texts, and so on.
5 Results and Discussion
The main findings are outlined below, each supported by the figures and tables provided. These findings can be further verified by the figures in Appendix C.
5.1 Model size: more parameters do not generally translate into better performance
The experiments conducted for text classification yield findings that challenge the assertion that larger models consistently deliver superior performance. An analysis of the impact of the logarithm of size on model performance indicates insufficient evidence to conclude that the Pearson correlation between size and performance is statistically significant in all the cases (refer to Fig. 2). This observed discrepancy may be attributed to several factors, including the relatively small size of the models considered, with a number of parameters up to billions rather than tens or hundreds of billions.
Specifically, in contextual embedding similarity, the trend of performance improvement with increasing size is almost nonexistent. As depicted in Fig. 3, for instance, models such as SapBERT (m06) and BioLORD (m07-m09), which excel in this approach, outperform models even three times larger in both datasets. When solely the performance deltas associated with increasing the number of parameters for the same type of model are examined, they are both positive and negative. For the transcriptions dataset, as indicated by the lighter crosses in Fig. 3, BERT (m01) marginally outperforms BERT (m00), similar to the case of BiomedBERT-large (abstracts only) (m04) versus BioMedBERT (abstracts only) (m03) where positive deltas are also obtained. However, for the MS-CXR dataset, BERT (m00) surpasses BERT (m01) in all metrics, and BioMedBERT (abstracts only) (m03) outperforms BiomedBERT-large (abstracts only) (m04) in precision. Now, by comparing BioMedBERT (abstracts only) (m03) and BiomedBERT (abstracts + full text) (m02), performance gains are obtained when more training data are used.
In natural language inference, while the results align with the expectation that larger models tend to lead to better performance, more experiments are needed to draw a solid conclusion. In the MS-CXR dataset, the difference between the most effective NLI-DeBERTa model and the least effective RoBERTa-MNLI and BART Large-MNLI models ranges from across the metrics considered. These differences suggest that larger models among the models used always lead to performance improvement. However, reaching the same conclusion for the transcriptions dataset is not straightforward, given the results especially obtained for RoBERTa-MNLI (m12), which, as depicted in Fig. 3, is closer to the performance of NLI-DeBERTa (m11) than to that of BART Large-MNLI (m13).
When it comes to multiple-choice question-answering, the impact of the increase in size on performance is also unclear. While positive Pearson correlations are obtained, there is insufficient evidence to deem them statistically significant. When analyzing this impact within the T5 model family, as shown in Fig. 3, the effect is minimal or inconsistent when considering their non-instruction-tuned versions (m14-m17, m24-m25). Conversely, when examining the instruction-tuned versions of T5 (m18-m23), a consistent positive effect of size on performance is observed for the transcriptions dataset. Although a positive effect is also present for the MS-CXR dataset, it is only for T0 models (m22-m23).
Overall, the results of the text classification task suggest that increasing model size only sometimes leads to improved performance, whether comparing different models or evaluating the same models. Therefore, while model size may be a relevant factor in determining performance, training data and objectives may be more decisive. Expanding the sample size could be essential to validate this observation, considering a minimum of 30 or 35 models per approach. The previous insights can be verified in more detail in Figs. 13, 14, 15, 16, 17 and 18 in Appendix C.
In the context of the conditional text generation task, as shown in Fig. 4, increasing the size of the model contributes to higher performances, regardless of whether the BOS token is included. Also, an improvement in the performance of the second versions (m42, m44) versus the first versions (m41, m43) of OpenLLaMA is observed. Considering that the difference between these OpenLLaMA versions is the dataset used for pre-training, the conditional text generation task results do not contradict the above conclusion for text classification. Furthermore, our findings do not conflict with existing scaling laws, usually associated with LLMs. First, because we have no control over training, i.e., we cannot guarantee that the model is not exposed to variations other than the number of parameters (size). Second, Chinchilla’s law does not consider inter-model comparisons.
5.2 Model domain: more than a specialized domain; model architecture, training data, and training objective
Current medical datasets remain relatively small compared to those of the general domain, covering only a tiny region of the medical knowledge space [79]. Since some of the available domain-specific models are specialized by continuous pre-training on only one of these datasets, their generalization ability may be hampered, counter to the goal of creating well-generalizable domain-specific models and even achieving the ideal of artificial general intelligence (AGI). As discussed previously, the results suggest that architecture, training data, and training objectives are crucial in determining model generalization, possibly outweighing model size.
For the classification task through contextual embedding similarity, the effectiveness of domain specialization in improving performance is not evident, as displayed in Fig. 3. For the transcriptions dataset, refer to the lighter blue crosses in the figure. For example, Bio+Clinical BERT (m10) achieves lower results than expected, being around the middle of the performance ranking. Even in biomedical domain specialization, BERT (m00), a general domain model, outperforms some of the BiomedBERT models (m02-m04), challenging the superiority of domain-specific models. These observations are valid for both datasets. In particular, the latter observation also extends to BERT (m01).
Two comparisons arise when employing multiple-choice question-answering: T5 models (m14-m15) versus their clinical specialized versions (m24-m25) and Alpaca (m39) versus MedAlpaca (m52). However, the evidence supporting the hypothesis that domain specialization significantly improves performance is limited, as displayed in Fig. 3. Specifically, in the transcriptions dataset, differences in AUC scores between ClinicalT5 (m24-m25) and T5 (m14-m15) are and for the base and large sizes, respectively. This trend persists across all prompts. Similarly, the difference in AUC scores between MedAlpaca (m52) and Alpaca (m39) is , negatively affecting performance. In the MS-CXR dataset, the difference in F1-scores between ClinicalT5 (m24-m25) and T5 (m14-m15) is and for the base and large sizes, respectively. For MedAlpaca (m52) versus Alpaca (m39), the F1-score deteriorates even by . Thus, the evidence suggests that domain specialization provides either minimal or nonexistent benefits for this task.
Similar patterns to those discussed previously are identified in the text generation task, as observed in Fig. 4. Remarkably, the second version of LLaMA, LLaMA 2-7B (m38), stands out as the model with the highest predictive capacity for the dataset, with an average perplexity of . In contrast, BioGPT (m48) presents the greatest difficulty in comprehending the dataset, with an average perplexity of . These results are paradoxical, especially considering that BioGPT (m48) is domain-specialized while LLaMA 2-7B (m38) is not. A more detailed analysis by size reveals that
1.
for the medium size, BioGPT (m48) does not outperform any of the general domain models, whereas GPT-2-PubMed Medium (m46) does;
2.
for the large size, domain specialization proves beneficial; and
3.
for the XL and XXL sizes, neither Galactica (m50-m51) nor BioGPT-Large (m49) clearly outperforms general domain models.
Consequently, the only specialized models that demonstrate an advantage are GPT-2-PubMed (m46-m47).
Considering the insights discussed and the remarkable performance of SapBERT (m06) and BioLORD (m07-m09) models in the contextual embedding similarity approach, as well as Flan-T5 (m18-m21) and T0 (m22-m23) in the multiple-choice question-answering approach, the fundamental role played by training data, training objectives, and model architectures is underscored. Continual pre-training for named entity recognition or medical entity linkage using contrastive learning on UMLS data might contribute to the success of SapBERT and BioLORD models in the contextual embedding similarity task. On the other hand, employing instruction-tuned text-to-text models represents a compelling approach to achieving high performance in multiple-choice question-answering. Additionally, exploring not only the hypothesis but also the impact of domain instruction fine-tuning could provide further insights into model performance and effectiveness in specialized domains. Given that it is impossible to conclude the impact of the model domain on NLI, expanding the analysis to incorporate domain-specialized NLI models in biomedical and clinical domains could be valuable.
5.3 Prompting and instruction-tuned models: key in determining model performance
One of the central points of the study is to analyze the influence of prompts on the models and text classification approaches under investigation. As shown in Fig. 5, the results underscore the fundamental role of the prompt and its wording in the performance of the model, demonstrating both positive and negative impacts. Consequently, we advocate using prompts and advanced prompt engineering techniques to guide the model toward better results. Nevertheless, this process should not be limited to a single prompt due to the observed and well-known phenomenon of prompt brittleness [11]. As well, instruction fine-tuning proves advantageous for the models. Detailed results for each classification approach are presented below.
In contextual embedding similarity, using a prompt does not always confer benefits. For the transcripts dataset, the mean impact on the AUC score is points, with differences ranging from to points. Overall, using any of the proposed prompts improves performance for of the model + pooling strategy combinations. As for the BioLORD-PMB, BiomedBERT family, BERT, and SciBERT models, none of the suggested prompts led to AUC score improvements. Inversely, for the MS-CXR dataset, the prompt impact on performance is generally more positive, albeit with more variability. The mean impact, along with its range, is (, ) for accuracy, (, ) for F1-score, (, ) for precision, and (, ) for recall. Overall, employing any of the proposed prompts represents benefits for to of the model + pooling strategy combinations, depending on the metric. As for the BioMedBERT (abstracts only) and BiomedBERT-large (abstracts only), their performance scores, except for precision, improve with any of the proposed prompts, while for the BioLORD family and Bio+Clinical BERT, their scores are always hindered.
When examining the prompt impact on NLI, more consistent benefits are observed than for contextual embedding similarity. Using any of the proposed prompts yields improvements in the transcriptions dataset. The mean impact is for BART Large-MNLI, for NLI-DeBERTa, and for RoBERTa-MNLI, with larger models profiting the most from its usage. On its side, in the MS-CXR dataset, using a prompt only sometimes results in gains, particularly for NLI-DeBERTa. For this model, the impact on the F1-score is , while for BART Large-MNLI and RoBERTa-MNLI, it is and , respectively. For NLI-DeBERTa, positive prompt impacts are only observed for precision. For NLI, there are certain prompts with a high positive impact, whereas others do not, mostly independent of the model.
Regarding multiple-choice question-answering, the importance of prompts is evident, as they can significantly influence model performance. For example, in the transcriptions dataset, the proportion of models performing better than random guessing (AUC scores greater than ) can improve from to with appropriate prompts. Similarly, in the MS-CXR dataset, this proportion can change from to when measured by the F1-score (scores greater than ). Thus, its importance is highlighted not only by the high performance obtained but also by its variability, as reflected in Fig. 5. This is further supported by Figs. 17 and 18 in Appendix C. When analyzing the impact of templates versus questions, both components of the prompt (see Appendix B), it is found that the wording of templates has more influence on performance than the wording of questions, and no single prompt (template + question) works universally well for all models.
More specifically, in the transcriptions dataset, the performance ranking between the T5 models and between the Flan-T5 and T0 models remains consistent independent of the prompt. Moreover, Flan-T5-XXL (m21) and T0++ (m23) consistently outperform the other models regardless of the prompt. On average, the impact of the template, measured by the standard deviation of the AUC score per model, is , while the effect of the question within a template is . For the MS-CXR dataset, it becomes evident that some models, such as T5 and Falcon-7B, are unsuitable for this approach and dataset. The prompt plays an even more critical role than in the transcriptions dataset, affecting the ranking of models. The sensitivity is such that, with certain prompts, instruction-tuned models yield similar results to their base counterparts. On average, the impact of the template is , while the effect of the question within a template is . Interestingly, one of the proposed role-based prompts works remarkably well for the instruction-tuned versions of the T5 family despite not being the optimal choice overall.
Concerning instruction-tuning, these models generally outperform their non instruction-tuned counterparts. The instruction-tuned T5 versions, whether T0 or Flan-T5, in any size considered, exhibit superior performance than their base counterparts. In the transcriptions dataset, Flan-T5 either performs better or is on par with T0. However, in the MS-CXR dataset, T0 outperforms Flan-T5 in the 11B parameter version only when comparing the highest performances due to the considerable variability of T0++. Also, instruction-tuning improves performance consistently for the LLaMA group, whereas this is not always the case for the other generative models: MPT and GPT-J are exceptions in the transcriptions dataset and Falcon in the MS-CXR dataset. Overall, instruction-tuning represents a gain, with mean improvements of in AUC-score and in F1 score between the highest performances per model and in AUC-score and in F1 score between the mean performances for the transcriptions dataset and the MS-CXR dataset, respectively.
More details on the prompt impact can be found in Figs. 20 and 20 in Appendix C.
5.4 Text classification: instruction-tuned T5: models to go for
| CES | NLI | MCQA | ||||||||
| Dataset | Metric | Model (poolingprompt) | Score | Model (prompt) | Score | Model (prompt) | Score | |||
| Accuracy | BioLORD-STAMB2-v1-STS2 | (avg5) | 69.65 | RoBERTa-MNLI | (5) | 76.50 | T0++ | (t03+q04, t03+q05) | 81.68 | |
| F1-score | BioLORD-STAMB2-v1-STS2 | (avg ) | 69.34 | RoBERTa-MNLI | (5) | 78.25 | T0++ | (t03+q05) | 83.90 | |
| MS-CXR | Precision | BioLORD-PMB | (CLS ) | 83.34 | RoBERTa-MNLI | (5) | 80.74 | Alpaca 7B | (t02+q03) | 85.85 |
| Recall | BioLORD-STAMB2-v1-STS2 | (avg ) | 72.58 | RoBERTa-MNLI | (5) | 82.28 | T0++ | (t03+q04) | 89.18 | |
| Transcriptions | AUC score | BioLORD-STAMB2-v1-STS2 | (avg1) | 89.01 | BART Large-MNLI | (4) | 80.75 | Flan-T5-XXL | (t01+q09) | 92.38 |
As shown in Table 2 and Fig. 3, for both datasets, the most favorable results are obtained with larger instruction-tuned T5 models, namely Flan-T5 (m18-m21) and T0 (m22-m23). The latter highlights the possibility of representing discriminative tasks as generative ones by framing them as instructions. It is also emphasized that generative tasks are not exclusive to decoder-only models, and text-to-text models might be an architecture that needs to be explored more and potentially promising. For example, instruction-tuned versions of T5 with 3B parameters (m20, m22) provide higher results than decoder-only models, almost three times larger, applying to both datasets.
When explicitly comparing precision and recall values, which can be critical depending on the application, the optimal choice of models may vary. Alpaca (m39) and LLaMA 2-CHAT-7B (m40) demonstrate proficiency in precision, while Flan-T5-XXL (m21) and T0++ (m23) excel in recall. Also, BioLORD (m07-m09) models, with 110 million parameters, exhibit precision and AUC score differences equal to or less than 3 points compared to models with a number of parameters between 7 billion and 11 billion.
With respect to the datasets, the MS-CXR dataset poses more challenges for the models than the transcriptions dataset. This difference in difficulty could be attributed to factors such as (i) the format of the data, (ii) the fact that the reports are from different medical subfields, and (iii) the length of the reports. Nevertheless, the highest performers per approach consistently show results indicative of clinical knowledge or clinical notions.
5.5 The groups of generative models: LLaMA and GPT
Two procedures are carried out to gain insight into whether the models exhibit similar perplexity behavior and to identify potential clusters among them. The first procedure involves calculating the correlations between the models. Spearman’s and Pearson’s correlations are considered, assessing monotonic and linear relationships. The second procedure involves dimensionality reduction via UMAP, followed by hierarchical clustering, represented in Fig. 6 using dendrograms.
The correlation analysis reveals that, in general, all models are positively related. This positive relation indicates that most samples have a similar relative difficulty level for the models. However, the BioGPT models are the only exceptions, showing low Spearman correlations compared to the other models and positioning them considerably far from the rest of the evaluated models. The observed behavior in terms of monotonic relationships remains consistent regardless of the use of the BOS token. Regarding linear relationships, groups of models are clearly observed, including LLaMA models, OpenLLaMA models, and GPT-2 models. Although the previous groups are somewhat expected, some unexpected associations are also evident, such as between Falcon-7B and MPT-7B and between Palmyra Base 5B and GPT-J 6B. These two groups have notably weak intergroup correlations.
With respect to the LLaMA and OpenLLaMA models, their linear relationship weakens when the BOS token is used, indicating more pronounced performance disparities. Possibly, training data plays an important role in this context, as it is essentially their main difference [92]. Further insights, provided by the correlations and supported by the dendrograms in Fig. 6, reveal the existence of two main groups of models: the GPT-2 models and the LLaMA models.
The analysis of the standard deviations reveals the presence of exceptionally challenging samples for the models, that is, outliers. This phenomenon is visually depicted in Fig. 4, which illustrates the distribution of perplexity between different models. Overall, the perplexity distributions have considerable variance, with the LLaMA models demonstrating the slightest variation. More detailed analyses are required to estimate the importance of these outliers and possible patterns among them.
6 Conclusion
This study comprehensively explored a range of language models with varying sizes and domain fields. These models, encompassing 52 in total, were tested in two fundamental natural language processing tasks: text classification and text generation. The size range of these models goes from 110 million to 13 billion parameters, which is relatively small compared to recent language models. Our findings, based on these specific models, have significant implications, particularly for researchers and organizations operating under resource-constrained contexts.
In text classification, our research has explored the effectiveness of three distinct approaches: context embedding similarity, natural language inference, and multiple-choice question-answering. BioLORD and SapBERT models have demonstrated remarkable performance in text classification via contextual embedding similarity. Additionally, the instruction-tuned versions of T5, Flan-T5 and T0, followed by the instruction-tuned versions of LLaMA, have exhibited good results in the multiple-choice question-answering approach. However, for the NLI approach, we have identified the need for extended exploration, particularly in the context of specialized domain models. This implies that further research is needed to fully understand the potential of NLI in text classification, especially in specialized domains.
A common thread running through our findings is the significance of the prompt in improving text classification performance across different datasets and approaches. The significance of prompts extends beyond mere performance gains; they present a viable alternative to the resource-intensive processes of training and fine-tuning language models, which are often associated with substantial financial and environmental costs, especially for large-scale models. Moreover, effective prompt engineering is essential to mitigate prompt brittleness, ensuring more robust and reliable outcomes.
Since, in general, medical datasets remain relatively small and cover only a small region of the medical knowledge space [79], domain-specific models, specialized by continuous pre-training on only one of these datasets, might see their generalization ability hindered. This practice could be one of the explanations for the results obtained. The results also suggest that the architecture, training data, and training objectives are crucial in determining the model’s generalization abilities, possibly outweighing model size.
Regarding the text generation task, two groups of models are identified based on the perplexities obtained on MIMIC-CXR. These groups consist of the GPT-2 models and the LLaMA models. Notably, the LLaMA models stand out due to their low perplexities with minimal variation. Further research is needed to identify and understand the outliers within these results, as they could hold important insights.
In conclusion, this research highlights the critical role of prompts in language model inference and reaffirms the effectiveness of instruction-tuned generative models in addressing downstream tasks. It also highlights the significance of model architecture, training data, and training objectives, potentially even more so than model size, in its generalization capacity. We advocate for further investigations into topics such as model calibration, i.e., how certain the model is about output, prompt tuning, and performance concerning issues like hallucinations and biases, among others. Such studies can lead to more effective and ethical applications of language models, particularly in sensitive fields like healthcare.
Appendix A Data
This section describes the data employed and outlines the corresponding preprocessing procedure.
A.1 Transcriptions
Transcriptions is a multi-label dataset with 40 different labels and data samples. The data were extracted from Kaggle, and additional information about the labels can be found in MTSamples.com.
A.1.1 Preprocessing
The preprocessing procedure involves the removal of samples that lack associated reports, adjusting the formatting of the report, and selecting and renaming labels. Formatting adjustments are necessary because line breaks are encoded as comma patterns. To ascertain the final format, we considered the original data source MTSamples.com and the results generated by ChatGPT as a guide to knowledge of language models.
In terms of labels, less relevant categories were excluded due to their broad level of generality or lack of association with a specific medical specialty. Precisely, the eliminated labels are: “Consult - History and Phy.”, “Discharge Summary”, “Emergency Room Reports”, “General Medicine”, “Hospice - Palliative Care”, “IME-QME-Work Comp etc.”, “Letters”, “Office Notes”, “Pain Management”, “SOAP / Chart / Progress Notes”. Additionally, several labels contained the “/” character, indicating “or”, which we explicitly replaced with the latter. For example, “Allergy / Immunology” was transformed into “Allergy or Immunology”. Subsequently, the labels “Chiropractic” and “Physical Medicine - Rehab” were merged into a unified category called “Physical Medicine and Rehabilitation, or Chiropractic”. Other modifications include transforming “ENT - Otolaryngology” into “Otolaryngology”, “Hematology - Oncology” into “Hematology or Oncology”, “Lab Medicine - Pathology” into “Laboratory Medicine or Clinical Pathology”, “Pediatrics - Neonatal” into “Pediatrics or Neonatal”, and “Speech - Language” into “Speech and Language”.
Upon completion of the preprocessing, the initial count of 40 different labels is reduced to 29, and the number of samples to consider is .
A.1.2 Description
The class distribution is visualized in Fig. 8. Surgery is the most prevalent category (in of the samples), followed by Cardiovascular or Pulmonary () and Orthopedic (). On the other hand, Allergy or Immunology (), preceded by Autopsy () and Laboratory Medicine or Clinical Pathology (), are the least frequent categories. The number of labels per sample ranges from 1 to 4, with an average of 2 labels per sample. Additionally, some labels never co-occur within the same sample.
The results of analyzing label leakage, which refers to whether a label appears explicitly in the text to be classified, are shown in Fig. 8. For most labels, label leakage is minimal, except for Autopsy (), Rheumatology (), Speech and Language (), and Surgery (). Labels without label leakage are Allergy or Immunology, Cardiovascular or Pulmonary, Cosmetic or Plastic Surgery, Diets and Nutritions, Hematology or Oncology, Laboratory Medicine or Clinical Pathology, Obstetrics or Gynecology, Pediatrics or Neonatal, Physical Medicine and Rehabilitation, or Chiropractic, Psychiatry or Psychology, and Sleep Medicine. The presence of labels in texts of other labels is not considered, given that this is a multi-label dataset, and the analysis and interpretation of such occurrences are inherently complex.
A.2 MS-CXR
MS-CXR [100, 101, 102] is a multi-class dataset with 8 different classes and a corpus of data samples, comprising 718 unique samples. The data can be obtained from [102].
A.2.1 Preprocessing
The preprocessing procedure involves removing instances without associated reports and eliminating duplicates. To be precise, 730 samples () were identified as duplicates, with a maximum of 82 and an average of 3 duplicates, considering only repeated reports. In addition, when duplicate reports do not agree with the assigned label, either of these labels is evaluated as the true one.
A.2.2 Description
The class distribution is depicted in Fig. 10. Overall, the dataset does not exhibit class imbalance. The most frequent classes are Pneumonia (), closely followed by Pneumothorax (), while the less frequent classes are Cardiomegaly (), preceded by Edema ().
Upon analysis of label leakage, as presented in Fig. 10, a high label leakage is observed, except for Lung Opacity, which has a low leakage rate of . In particular, Consolidation, Edema, and Pneumothorax exhibit leakage rates that exceed . Classes with leakage rates below include Pneumonia, Cardiomegaly, and Lung Opacity, as mentioned earlier. Regarding the presence of labels in text from other labels, notable occurrences include Consolidation in the classes of Edema () and Pneumonia (), and Pleural Effusion in the Atelectasis class ().
To conclude, each class’s word count per text is measured, and their distributions are presented in Fig. 12. Classes with shorter texts include Cardiomegaly and Pneumothorax. Although classes with longer texts are not observed, there are flatter distributions with heavy tails, suggesting that the length of texts in these classes is less concentrated around a specific value.
A.3 MIMIC-CXR
MIMIC-CXR [103, 104, 102] is a dataset of radiographic reports that encompasses samples. After extracting the most pertinent sections, samples are identified as informative. This dataset is accessible through [102].
A.3.1 Preprocessing
The preprocessing procedure involves extracting the most relevant sections from chest X-ray reports using the codes [105] designed for this purpose and publicly available on GitHub [106]. In addition, texts lacking content and duplicate samples are removed. Texts lacking information are defined as those that are empty or match one of the following: “.”, “As above”, “As above.”, “As above..”, “None.”, “See above.”, “No changes.”, “___”, “___ earlier”, “___,”, or “___.”. Those mentioned above were identified after meticulously examining texts with a maximum length of two words. In total, these non-informative texts represent merely of the dataset. Regarding duplicates, of the total samples are duplicated, comprising of the dataset. The text with the most duplicates is ”No acute cardiopulmonary process.” representing of the samples. On average, each text appears twice in the dataset.
Upon completion of the preprocessing steps, the dataset results in , composed mainly of impressions () and findings ().
A.3.2 Description
Considering the nature of this dataset, its description focuses mainly on the distribution of the number of words per sample, as shown in Fig. 12. This distribution is left-skewed, with a peak of around 10 words per sample. Moreover, there is a significant plateau between 20 and 40 words per sample. Interestingly, the distribution’s right tail extends beyond 150 words per sample. In summary, most texts () contain at most 51 words, with a pronounced peak of around 10 words per sample. However, this dataset also includes longer texts, some reaching up to 307 words.
Appendix B Prompts
The prompts used for the text classification task via contextual embedding similarity, natural language inference (NLI), and multiple-choice question answering (QA) are presented.
B.1 Prompts for text classification via contextual embedding similarity and NLI
The prompts proposed for text classification using contextual embedding similarity and Natural Language Inference (NLI) are exclusively applied to the label (in the case of NLI, to the hypothesis). Table 3 lists the prompts used. Template ID 0 is the default to generate the hypothesis in the zero-shot text classification using the NLI setting, as documented in HuggingFace.
| ID | Label prompt | Dataset |
| 0 | This example is {label}. | Transcriptions, MS-CXR |
| 1 | This is an example of {label}. | Transcriptions, MS-CXR |
| 2 | This report belongs to the category {label}. | Transcriptions |
| 3 | This report belongs to the medical speciality {label}. | Transcriptions |
| 4 | This report belongs to the medical speciality: {label}. | Transcriptions |
| 5 | The diagnosis is {label}. | MS-CXR |
| 6 | There is evidence of {label}. | MS-CXR |
| 7 | These findings are consistent with {label}. | MS-CXR |
B.2 Prompts for text classification via multiple-choice QA
The proposed prompts for text classification via multiple-choice question answering are based on the default templates specific to various of the considered instruction-tuned models. These templates are systematically assessed using a set of questions, enabling us to quantify the influence of the question wording. For the MS-CXR dataset, we also incorporate role-based questions. The prompts, their corresponding datasets, and specific requirements are summarized in Table 5 and Table 4.
Each class or label is encoded with an uppercase letter denoting the option, followed by its name. For instance, if the first label is “”, it is represented as “(A) ” within the prompt. In the context of the transcriptions dataset, there are 29 distinct labels. However, due to their large number, we include the top 10 most frequent labels and categorize the remaining labels under an additional “Other” option. Specifically, for the transcriptions dataset, we employ templates t01, t02, t03, t04, and t07 along with questions q07, q08, and q09. Whereas for MS-CXR dataset, we employ templates t01, t02, t03, t04, t07, t11, and t13, and questions q03, q04, and q05.
| ID | Question | Dataset |
| q01 | What is the most plausible diagnosis? | MS-CXR |
| q02 | What is the patient’s diagnosis? | MS-CXR |
| q03 | What is the diagnosis? | MS-CXR |
| q04 | Which one of the following is the diagnosis? | MS-CXR |
| q05 | Which one is the patient’s diagnosis? | MS-CXR |
| q06 | Which of the options is the most likely to be the diagnosis? | MS-CXR |
| q07 | Which category does the report belong to? | Transcriptions |
| q08 | What is the field that best suits the report? | Transcriptions |
| q09 | Which one is the topic of the report? | Transcriptions |
| ID | Prompt | Requirements | Dataset |
| t01 | Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request. ### Instruction: {question} Select one of the following options: {options} ### Input: {report} ### Response: ( | report, options, QUESTION | Transcriptions, MS-CXR |
| t02 | Context: {report} Question: {question} Options: {options} Answer: ( | report, options, QUESTION | Transcriptions, MS-CXR |
| t03 | Context: {report} Question: Based on the context, {question} Options: {options} Answer: ( | report, options, question | Transcriptions, MS-CXR |
| t04 | {report}. Which one of the following, if true, most strengthens the argument? {options}. ( | report, options | Transcriptions, MS-CXR |
| t05 | Read the following and answer the question. {report} {question} {options} ( | report, options, QUESTION | Transcriptions, MS-CXR |
| t06 | {report} What’s the best answer to this question: {question} {options} ( | report, options, QUESTION | Transcriptions, MS-CXR |
| t07 | {report} {question} {options} ( | report, options, QUESTION | Transcriptions, MS-CXR |
| t08 | Read this chest x-ray report: “{report}” Now answer this question: “{question}” {options} ( | report, options, QUESTION | MS-CXR |
| t09 | Knowing that “{report}”, how would one answer “{question}” {options} ( | report, options, QUESTION | Transcriptions, MS-CXR |
| t10 | {report} Based on the above text, what’s the best answer to this question: {question} {options} ( | report, options, QUESTION | Transcriptions, MS-CXR |
| t11 | You are a doctor and have the following information about a patient from a chest x-ray: {report}. Which one of the following, if true, most strengthens the argument? {options}. ( | report, options | MS-CXR |
| t12 | You are a doctor and have the following information about a patient from a chest x-ray: {report}. {question} {options}. ( | report, options, QUESTION | MS-CXR |
| t13 | I want you to act as a virtual doctor. I will describe my symptoms and you will choose the most probable diagnosis among the following: {options}. You should only reply with the chosen diagnosis, and nothing else. My request is “{report}”. ( | report, options | MS-CXR |
| t14 | I want you to act as a virtual doctor. I will describe my symptoms and you will choose a diagnosis among the possible diag- noses. You should only reply with the chosen diagnosis, and nothing else. Do not write explanations. The possible diagnoses are: {options}. My request is “{report}”. ( | report, options | MS-CXR |
Appendix C Supplementary results
This appendix presents supplementary figures and tables to support the results presented. They are displayed first by dataset and then by task or approach.
C.1 Classification task
C.1.1 Contextual embedding similarity
Results are depicted in Figs. 3, 14 and 13. The mapping between the models and their ID is
| m00: BERT | m06: SapBERT |
| m01: BERT | m07: BioLORD-STAMB2-v1 |
| m02: BiomedBERT (abstracts + full text) | m08: BioLORD-STAMB2-v1-STS2 |
| m03: BiomedBERT (abstracts only) | m09: BioLORD-PMB |
| m04: BiomedBERT-large (abstracts only) | m10: Bio+Clinical BERT |
| m05: SciBERT |
C.1.2 Natural language inference
Results are depicted in Figs. 16 and 15. The mapping between the models and their ID is
| m11: NLI-DeBERTa | m12: RoBERTa-MNLI | m13: BART Large-MNLI |
C.1.3 Multiple choice question answering
| m14: T5-V1.1-Base | m21: Flan-T5-XXL | m31: GPT-J 6B | m38: LLaMA 2-7B |
| m15: T5-V1.1-Large | m22: T0-3B | m32: Instruct GPT-J | m39: Alpaca 7B |
| m16: T5-V1.1-3B | m23: T0++ | m33: Falcon-7B | m40: LLaMA 2-CHAT-7B |
| m17: T5-V1.1-11B | m24: ClinicalT5-base | m34: Falcon-7B-Instruct | m52: MedAlpaca 7b |
| m18: Flan-T5-Base | m25: ClinicalT5-large | m35: MPT-7B | |
| m19: Flan-T5-Large | m29: Palmyra Base 5B | m36: MPT-7B-Instruct | |
| m20: Flan-T5-XL | m30: Camel 5B | m37: LLaMA-7B |

### v2

@@
 Abstract
-Since the emergence of the Transformer architecture, language model development has increased, driven by their promising potential.
-However, releasing these models into production requires properly understanding their behavior, particularly in sensitive domains such as medicine.
-Despite this need, the medical literature still lacks technical assessments of pre-trained language models, which are especially valuable in resource-constrained settings in terms of computational power or limited budget.
-To address this gap, we provide a comprehensive survey of language models in the medical domain.
-In addition, we selected a subset of these models for thorough evaluation, focusing on classification and text generation tasks.
-Our subset encompasses 53 models, ranging from 110 million to 13 billion parameters, spanning the three families of Transformer-based models and from diverse knowledge domains.
-This study employs a series of approaches for text classification together with zero-shot prompting instead of model training or fine-tuning, which closely resembles the limited resource setting in which many users of language models find themselves.
-Encouragingly, our findings reveal remarkable performance across various tasks and datasets, underscoring the latent potential of certain models to contain medical knowledge, even without domain specialization.
-Consequently, our study advocates for further exploration of model applications in medical contexts, particularly in resource-constrained settings.
+Since the Transformer architecture emerged, language model development has grown, driven by their promising potential.
+Releasing these models into production requires properly understanding their behavior, particularly in sensitive domains like medicine.
+Despite this need, the medical literature still lacks practical assessment of pre-trained language models, which are especially valuable in settings where only consumer-grade computational resources are available.
+To address this gap, we have conducted a comprehensive survey of language models in the medical field and evaluated a subset of these for medical text classification and conditional text generation.
+The subset includes models with million to billion parameters, spanning the Transformer-based model families and knowledge domains.
+Different approaches are employed for text classification, including zero-shot learning, enabling tuning without the need to train the model.
+These approaches are helpful in our target settings, where many users of language models find themselves.
+The results reveal remarkable performance across the tasks and datasets evaluated, underscoring the potential of certain models to contain medical knowledge, even without domain specialization.
+This study thus advocates for further exploration of model applications in medical contexts, particularly in computational resource-constrained settings, to benefit a wide range of users.
 The code is available on https://github.com/anpoc/Language-models-in-medicine.
 1 Introduction
-Natural Language Processing (NLP) holds immense promise in the medical domain.
-Recently, the medical community has shown substantial interest in leveraging state-of-the-art language models to address various medical challenges [1, 2].
-In particular, generative large language models (LLMs) have showcased emergent abilities beyond their original training objectives, such as text summarization and question-answering [3].
+Natural language processing (NLP) holds great promise in the medical field.
+The medical community has recently shown substantial interest in leveraging state-of-the-art language models to address various medical challenges [1, 2].
+In particular, generative large language models (LLMs) have showcased emergent abilities beyond their original training objectives, such as text summarization and question answering [3].
 These newfound abilities have enabled LLMs to perform tasks of significant clinical importance, including passing medical examinations, summarizing clinical and radiological reports, as well as medical dialogues, extracting drug names from medical notes, responding to patient inquiries, and writing medical histories and physical assessments [2, 4].
@@
 Integration of language models in medical settings is becoming a reality as partnerships between developers and healthcare systems continue to grow [7].
-The potential benefits are significant, as they can derive broadly applicable representations from extensive medical corpora at scale.
-Recent research indicates that LLMs with more than 100 billion parameters can encapsulate clinical knowledge [8].
-Nevertheless, it is essential to recognize that our understanding of the behavior of LLMs and language models in general still needs to be completed [4].
-Moreover, implementing these models also carries risks, such as the generation of inaccurate results, a phenomenon known as hallucinations, and the potential amplification of existing biases [1, 4].
-Hence, the deployment of language models in sensitive fields, such as healthcare, should be approached with the utmost care [5].
-To further understand and assess language models’ performance in clinical scenarios, we conducted comprehensive evaluations using models of varying sizes, ranging from 110M to 13B parameters, and encompassing different knowledge domains.
-We focus on text classification, using three different approaches and conditional text generation tasks.
-The approaches employed for text classification and the application of zero-shot prompting alongside allow performance improvement without the need for training or fine-tuning.
-For its part, the selection of the model size range considers resource constraints, in terms of limited budgets or computational power, encountered not only by many research groups but also by other actors and institutions, such as hospitals, where handling models with more than 110 billion parameters is not feasible.
-Moreover, the use of small language models aligns with the crucial imperative of environmental sustainability, as larger models often require significant energy and computational resources.
-It also opens up possibilities for organizations to develop applications that can run directly on commodity hardware and small devices rather than relying on cloud-based services.
-In fact, the prospect of deploying these language models on smartphones and other mobile devices operating at the edge without cloud dependence is promising [9].
-Therefore, the use of language models in resource-constrained settings not only addresses practical challenges but also has great potential in edge and local computing.
-The paper is structured as follows.
-Section 1 introduces the study and provides an overview of its objectives.
-Section 2 elaborates on the preliminaries, establishing foundational concepts relevant to the investigation.
-Related work in the existing literature is explored in Section 3.
-Section 4 details the methodology adopted, and Section 5 provides the most significant results and analysis.
-Finally, Section 6 contains the conclusions drawn from the study, summarizing key insights and implications for future research directions.
+The potential benefits are significant, as they can derive broadly applicable representations from extensive medical corpora at scale and encapsulate clinical knowledge [8].
+Nevertheless, it is essential to recognize that our understanding of the behavior of both small pre-trained and large language models still needs to be completed [4].
+Deploying these models also carries risks, such as the generation of inaccurate results, a phenomenon known as hallucinations, and the potential amplification of existing biases [1, 4].
+Language models’ implementation in sensitive fields, such as healthcare, should therefore be approached with the utmost care [5].
+Computing and energy resources required by language models for their development and operation are another critical and limiting factor, especially in LLMs.
+The standard computing resources available in hospitals are of the consumer-grade type, where it is currently infeasible to handle models with hundreds of billions of parameters.
+Such resource-constrained settings, i.e., with consumer-grade computing resources, are presented not only by healthcare agents and institutions but also by research groups.
+When large language models do not represent a cost-effective or viable solution, smaller pre-trained language models can be an alternative.
+LLMs, albeit more massive, have similar architectures and pre-training tasks to smaller pre-trained language models [9].
+With the same computing budget, a smaller model trained with more high-quality data can perform better than its larger counterparts due to undertraining [10].
+Using curated scientific and biomedical corpora in pre-trained language models has also been effective for discriminative and generative language modeling [11].
+Furthermore, these smaller models align with the crucial imperative of environmental sustainability and open up the possibilities for organizations to develop applications that can run directly on commodity hardware and small devices rather than relying on cloud-based services [12].
+Language models in resource-constrained settings thereby address practical challenges and have great potential in local computing.
+To further understand the performance of language models in clinical scenarios with limited computational resources, we conducted a comprehensive evaluation focusing on the classification and conditional generation of medical texts in open-source models.
+The datasets employed enable the assessment of general and radiology-specific medical knowledge.
+In total, models are tested, ranging from M to B parameters, spanning all Transformer-based model families and knowledge domains from general to clinical.
+For conditional text generation, solely decoder-only models are used.
+The approaches adopted for text classification, together with prompt engineering, allow for improved model performance without the need for training or fine-tuning.
+An analysis of the impact of the prompts on performance is also included.
+To the best of our knowledge, this is the first work to evaluate such a large number of small pre-trained language models for medical tasks.
 2 Preliminaries
@@
 Each of these groups represents a paradigm shift in natural language modeling and has contributed significantly to the conception of language models as we know them today.
-The first transition, from statistical to neural language models, marked a shift from predicting words based on minimal local context to probabilistically assessing word sequences using neural networks.
-Additionally, this transition introduced the concept of representing words as low-dimensional continuous embeddings based on their contextual usage (distributional semantics).
-The second transition, from neural language models to pre-trained language models, represented a move from task-specific model development towards pre-training and fine-tuning methodologies.
-Lastly, the third transition to large language models constituted a transformation from discriminative AI to generative AI, from model-centric to data-centric approaches, and from fine-tuning to prompt-tuning [10, 11, 12].
+The first transition, from statistical to neural language models, entailed a shift from word prediction based on minimal local context to probabilistic evaluation of word sequences using neural networks.
+This transition also introduced the representation of words as low-dimensional continuous embeddings based on their contextual usage (distributional semantics).
+The second transition, from neural to pre-trained language models, involved turning from developing task-specific models to pre-training and fine-tuning methodologies.
+The third transition to large language models moved the focus from discriminative AI to generative AI, from model-centric to data-centric approaches, and from fine-tuning to prompt engineering and prompt tuning [9, 13, 14].
 These advances have paved the way for more sophisticated language models with broader applications and improved capabilities.
-2.1 From task-specific model development towards pre-training and fine-tuning: pre-trained language models
-The emergence of pre-trained language models gave relevance to pre-training and fine-tuning methodologies, which was a paradigm shift, as mentioned above.
-This shift has driven research toward designing more efficient architectures and refinement of pre-training strategies, demonstrating significant advantages in enriching language understanding and improving model performance in various tasks [11, 10].
-One of the first attempts at pre-trained language models is ELMo [13].
-ELMo was developed to capture context-aware word representations by pre-training a bidirectional Long Short-Term Memory (biLSTM) network and fine-tuning it for subsequent downstream tasks.
-However, the introduction of the Transformer architecture represented a qualitative leap in NLP, offering highly parallelizable structures and self-attention mechanisms.
-The Transformer [6] follows the autoencoder archetype, giving rise to three families of models: (1) BERT-family or encoder-only models, (2) GPT-family or decoder-only models, and (3) text-to-text or encoder-decoder models.
-Their graphical representations are shown in Fig.
-1.
+2.1 Pre-trained language models
+The emergence of pre-trained language models represented a paradigm shift, driving research toward designing more efficient architectures and refining pre-training strategies.
+These pre-trained models have been commonly adapted or specialized to downstream tasks via fine-tuning, which involves transferring knowledge by further training a model on new data.
+There are significant advantages demonstrated by these models in language understanding and model performance in various tasks [13, 9].
+ELMo is one of the earliest attempts at pre-trained language models[15].
+This model was developed to capture context-aware word representations by pre-training a bidirectional Long Short-Term Memory (biLSTM) network and fine-tuning it for subsequent downstream tasks.
+Later the Transformer architecture was introduced, revolutionizing the NLP field by offering highly parallelizable structures and self-attention mechanisms.
+The Transformer [6] follows the autoencoder archetype, from which three families of models arose: (1) BERT-family or encoder-only models, (2) GPT-family or decoder-only models, and (3) text-to-text or encoder-decoder models.
+In Fig.
+1, the graphical representations of these families are shown.
 2.1.1 Encoder-only models
@@
 2.1.2 Decoder-only models
-Decoder-only models focus on autoregressive language modeling, that is, predicting the next token in a sequence based on previous tokens.
-These models produce contextual embeddings and a distribution over the next tokens , which can be stated as .
+Decoder-only models focus on autoregressive language modeling, i.e., predicting the next token in a sequence based on previous tokens.
+These models produce contextual embeddings and distribution over the subsequent tokens , which can be stated as .
 However, the contextual embeddings they generate depend solely on the left context.
-Currently, most research efforts are directed toward decoder-only models due to their exceptional performance in conditional generation tasks and their demonstrated emergent capabilities.
+Most research efforts are currently directed toward decoder-only models due to their exceptional performance in conditional generation tasks and their demonstrated emergent capabilities.
 2.1.3 Encoder-decoder models
@@
 These models combine bidirectional contextual embeddings with the capability to generate output sequences, making them versatile in various text-to-text tasks without requiring additional heads for fine-tuning.
-Moreover, by having a wide spectrum of language tasks that can be translated into text-to-text representation, these models can potentially be used for a wide range of applications.
-2.2 From discriminative AI to generative AI: large language models
-More recently, it has been observed that increasing the size of language models, whether in terms of model size, dataset size, or computational resources, often results in improved model capabilities in various tasks [15, 16, 17, 18, 19, 20, 21], including those that require specialized scientific knowledge and reasoning [22].
-Known as the scaling law [23], this phenomenon posits a power-law relationship between model performance, measured by the loss, and the factors mentioned above.
-Several studies have delved into this relationship by progressively training larger language models, up to hundreds of billion parameters, such as GPT-3 [17], PaLM [19], Galactica [24], LLaMA models [25, 26], Claude [27], Gemini 1.5 [28], and Mistral [29].
-Although scaling primarily involves increasing model size while retaining similar architectures and pre-training tasks as smaller models [10], large language models (LLMs) have exhibited unexpected emergent capabilities [30, 3], enabling them to address a wide range of complex tasks.
-GPT-3 and ChatGPT serve as precursors to these large language models [11, 10].
-A particularly intriguing aspect of LLMs is their few-shot and zero-shot capabilities [31], enabling them to adapt to various tasks without requiring gradient-based parameter updates [8, 12, 32, 17, 21, 33, 34, 35].
-Through these and other prompting techniques, emerging capabilities have been observed, enabling LLMs to exhibit apparent reasoning abilities [8].
-Nonetheless, prompting techniques are not exclusive to LLMs but are also applicable, for example, to pre-trained language models.
-Similarly, numerous studies highlight the competence of LLMs as implicit knowledge bases [8, 22, 24, 36].
-As for prompt engineering, its role is highly significant due to its benefits, which have proven to be more effective in large or instruction-tuned models.
-[21, 32, 8].
-In the medical domain, for example, Chain of Thought (CoT) has been used to address explainability [37] and in-context learning to mitigate the need for costly medical annotations [11].
-Among LLMs, GPT-4 stands out for its exceptional performance, often matching or surpassing human performance in various tasks [38, 39, 40], including in scientific domains such as biology, chemistry, and medicine [41].
-In the general domain, extensive evaluations of this model have been conducted [22, 42, 43, 44], even exploring the path toward Artificial General Intelligence (AGI) [39].
-In the medical domain, recent studies have begun to investigate its utility in various applications, ranging from medical chatbots [45] to medical competency exams [40] and its applications in radiology [46, 38, 47, 48, 49, 50, 51], among others [52, 53, 54, 55].
-Nevertheless, there are other relevant models in the medical field, and unlike GPT-4, most of them are domain-specific.
-Generally, these models are derived from general-purpose LLMs, with exceptions like GatorTron [56], Galactica, and GatorTronGPT [57].
-The most common training technique is instruction fine-tuning, as seen in Galactica, MedAlpaca [58], PMC-LLaMA [59], Med-PaLM 2 [60], GatorTronGPT, and ClinicalGPT [61].
-However, although less frequently, reinforcement learning from human feedback (RLHF) and reinforcement learning from AI feedback (RLAIF) techniques are also adopted, HuatuoGPT [62] being an example.
-Recent research studies also indicate a multimodal trend that supports various types of healthcare data, including electronic health records (EHR), medical images, and medical sequence signals.
-Examples of these developments include LLaVAMed [63], MedAGI [64], OphGLM [65], Visual Med-Alpaca [66], and MedFlamingo [67], CheXzero [68].
+Moreover, by having a broad spectrum of language tasks that can be translated into text-to-text representation, these models can potentially be used for a wide range of applications.
+2.2 Large language models
+Scaling of language models has often resulted in improved model capabilities in various tasks [17, 10, 18, 19, 20, 21, 22], including those requiring specialized scientific knowledge and reasoning [23].
+Research by Kaplan et al.
+[24] revealed that there is an empirical power-law relationship between the language model performance, in terms of cross-entropy loss, and the model size, dataset size, and amount of compute used for training.
+It was further found that architectural details, such as network width or depth, had minimal effects on performance.
+Scaling laws have further been studied by Hoffmann et al.
+[10] and Bahri et al.
+[25].
+Following these empirical results, several studies have trained progressively larger language models of up to hundreds of billion parameters, such as GPT-3 [18], PaLM [20], Galactica [26], LLaMA [27, 28], Claude [29], Gemini 1.5 [30], and Mistral [31].
+Among all, GPT-3 and ChatGPT can be considered the precursors of the large language models, the name by which these large-scale language models are known [13, 9].
+GPT-4, a latter version of GPT-3, stands out for its exceptional performance, often matching or surpassing human performance on a variety of tasks [11, 32, 33], even in specialized domains [34].
+Extensive evaluations have been conducted to GPT-4 [23, 35, 36, 37], exploring even the path toward Artificial General Intelligence (AGI) [32].
+LLMs can be adapted to different tasks via prompt engineering, which, unlike fine-tuning, does not require retraining the model and updating its weights.
+These prompting techniques have led to observing unexpected emergent capabilities in LLMs, demonstrating the potential to address a wide range of complex tasks and exhibit apparent reasoning abilities [38, 3, 8, 14, 39, 18, 22, 40, 41, 42].
+In the medical field, for example, Chain of Thought (CoT) has been used for explainability [43] and in-context learning to mitigate the need for costly medical annotations [13].
+Numerous studies have even highlighted the competence of large language models as implicit knowledge bases [8, 23, 26, 44].
+In-context learning techniques, such as zero-shot and few-shot learning, have also proven to be remarkably effective on instruction-tuned models and models to which reinforcement learning techniques have been applied [22, 39, 8, 45].
+Zero-shot learning consists of asking a trained model to complete a task without providing explicit examples of this task, whereas in few-shot learning, some examples are provided.
+Nonetheless, prompting techniques are not exclusive to LLMs but are also applicable to smaller pre-trained language models, especially encoder-decoder and decoder-only models.
 Despite their advantages, LLMs also have limitations.
-One of these are computational challenges due to their high computational resource requirements [8, 10].
-Conducting repetitive and ablation studies to investigate the impact of training strategies for LLMs is costly and time-consuming.
-Therefore, in certain cases, LLMs may not represent a cost-effective solution.
-Since smaller language models have also shown the effectiveness of using curated scientific and biomedical corpora for both discriminative and generative language modeling [38], they represent an alternative to LLMs.
-Such models include BioBERT [69], SciBERT [70], PubMedBERT [71], BioMegatron [72], ScholarBERT [73], BioGPT [74], and ClinicalBERT [75].
-However, the current literature still lacks comparative studies on these smaller language models.
+Their high computational resource requirements and associated computational challenges represent a major limitation of these models.
+For example, conducting studies to better understand the behavior of LLMs and assess important criteria, such as faithfulness and biases, can be costly and time-consuming.
+The detection biases and hallucinations, i.e., generating inaccurate results, is crucial in sensitive domains such as medicine.
+Due to the significance of the computing limitation, alternatives such as model quantization [46] have been introduced.
+Quantization is a technique that reduces the computational and memory costs of model inference by representing its weights and activations with low-precision data types, such as 8-bit integers, instead of the usual 32-bit floating point.
+In natural language processing, this technique is currently being extensively studied, with [47, 48, 49, 50, 51, 52] being some examples in the literature.
+Recommendations on the optimal use of computational resources have also been proposed.
+Chinchilla’s scaling law [10], one of these recommendations, states that the optimal model size and the number of tokens for training a language model should scale equally for compute-optimal training under a given computational budget.
+In [10], it is further proved that current large language models are significantly undertrained due to the recent focus on scaling language models while keeping the amount of training data constant.
+A smaller model trained with more high-quality data can thus achieve better performance than its larger counterparts with the same computing budget.
+2.3 Language models in the biomedical/clinical context
+Broadly speaking, language models used in specialized domains are (i) trained models solely on target domain data, (ii) pre-trained models on general domain corpus with tuning strategies, and (iii) pre-trained models on specialized domain corpus with(out) tuning strategies.
+Examples of tuning strategies are fine-tuning with target domain data and prompt engineering.
+Pre-training on (ii) can also be domain-adaptive continual pre-training (i.e., pre-trained on specialized domain corpus after pre-training on general domain corpus) or mixed-domain pre-training (i.e., pre-trained on a mix of general and specialized domain corpus, simultaneously).
+GPT-4 is an example of a general domain language model that has been studied in medical applications.
+Research has covered from its utility as a medical chatbot [53] and in medical competency exams [33] to its applications in radiology [54, 11, 55, 56, 57, 58, 59], among others [60, 61, 62, 63].
+Nevertheless, the models studied in the medical context are mostly domain-specific, either biomedical or clinical.
+These models include pre-trained models such as BioBERT [64], SciBERT [65], BioMedBERT [66], BioMegatron [67], ScholarBERT [68], BioGPT [69], and ClinicalBERT [70]; as well as large language models as Galactica [26], MedAlpaca [71], PMC-LLaMA [72], Med-PaLM 2 [73], GatorTron [74], GatorTronGPT [75], and ClinicalGPT [76].
+Domain-specific models usually contain general domain data within the pre-training data, with exceptions such as BioMedBERT, Galactica, GatorTron, and GatorTronGPT.
+For large language models, instruction fine-tuning is the most common tuning technique, as in MedAlpaca, Med-PaLM 2, GatorTronGPT, and ClinicalGPT.
+Reinforcement learning from human feedback (RLHF) and reinforcement learning from AI feedback (RLAIF) have also been adopted, although less frequently, being HuatuoGPT [77] an example of this.
+Recent research studies indicate as well a multimodal trend that supports various types of healthcare data, including electronic health records (EHR), medical images, and medical sequence signals.
+Examples of these developments include LLaVAMed [78], MedAGI [79], OphGLM [80], Visual Med-Alpaca [81], MedFlamingo [82], and CheXzero [83].
 3 Related Work
-Comparative studies investigating language models are crucial in advancing our understanding of them, shedding light on their functionalities, and pinpointing their constraints.
-Despite previous research, a notable gap persists in the literature.
-This gap is particularly significant in fields that require heightened sensitivity, such as medicine, where a thorough understanding of models is imperative.
-Although existing research in this domain focuses primarily on specific tasks, datasets, or models [5, 32, 12, 38, 8], there are exceptions.
-Some of these studies are detailed below.
-In [11], the authors conduct a descriptive study delineating the capabilities and limitations of language models within the healthcare domain, encompassing pre-trained and large language models.
-They explain the development process, provide insight into training data, methodologies, and optimization strategies, and investigate concerns about integrating LLMs in healthcare.
-These concerns focus on fairness, accountability, transparency, and ethics.
-Likewise, in the context of integration concerns, the faithfulness problem within existing research on artificial intelligence in healthcare and medicine is explored in [76].
-This analysis involves identifying the causes of nonfaithful results, possible evaluation metrics for faithfulness, and potential mitigation strategies for this issue.
-Regarding studies closely aligned with the present work, notable contributions include those by Soni et al.
-[77], Lehman et al.
-[78], and Zhou et al.
-[79].
-In [77], the authors assess the cost-effectiveness of pre-training and fine-tuning in Transformer language models, specifically in BERT, BioBERT [69], Clinical BERT [75], and XLNet [80].
-These models underwent pre-training and fine-tuning on general, biomedical, and clinical-domain data combinations for medical question-answering (QA) tasks.
-Their findings revealed that BERT-based models exhibit superior performance when fine-tuned on multiple datasets, incorporating both general and clinical domain data.
-This observation suggests a gap in well-generalizable medical QA datasets.
-Furthermore, their results underscored the positive impact of initial fine-tuning on general-domain datasets, such as SQuAD, before fine-tuning on clinical datasets, highlighting the quality of the large, manually constructed SQuAD dataset.
-Overall, it was observed that initial fine-tuning improves performance in most cases.
-It is important to note that this study focused solely on question-answering and did not evaluate prompting techniques.
-On its side, in [78], the authors’ research delves into whether LLMs, primarily trained on general web text, are suitable for highly specialized, safety-critical domains such as medicine or if domain-specialized models offer better performance.
-The authors evaluated 12 language models, ranging from 220 million to 175 billion parameters, measuring their performance in three clinical tasks.
-As part of their experiments, T5 models were trained from scratch using MIMIC-III and MIMIC-IV clinical notes to investigate the efficiency of clinical tokens.
-Their findings suggest that relatively small specialized clinical models significantly outperform all in-context learning approaches, even when fine-tuned on limited annotated data.
-However, this study does not assess the ability of models to handle long text and considered neither decoder-only models nor instruction-tuned ones.
-Lastly, [79] provides a comprehensive overview of the development and deployment of LLMs in medicine, including the challenges and opportunities that they face.
-The main highlight of this study is that it is not only a discursive but also a practical assessment.
-The authors begin with an introduction to the principles of existing medical LLMs, including their basic model structures, the number of parameters, and the sources and scales of data used for model development.
-A comparative evaluation of the performance of different LLMs in various medical tasks is then provided, comparing them with state-of-the-art lightweight models.
-This review is framed by the research questions: 1) What are the practices for developing medical LLMs?
-2) How to measure the medical task performance of LLMs in a medical setting?
-3) How have medical LLMs been used in actual practice?
-4) What are the challenges in using medical LLMs; and 5) How to develop and deploy medical LLMs more effectively?
+Comparative studies investigating language models are crucial to advance our understanding of them, shed light on their functionalities and pinpoint their constraints.
+Despite previous research, a notable gap persists in the literature due to, among other cause, current pace of development in NLP.
+This gap is particularly significant in fields that require heightened sensitivity, such as medicine, where a thorough understanding of models is imperative [45].
+Existing research in medicine is mainly focused on specific tasks or datasets or models [5, 39, 14, 11, 8].
+Moreover, most of the discursive and practical assessments focus on LLMs, as can be seen below.
+To the best of our knowledge, there is no practical assessment in the clinical context that includes a wide number of pre-trained models, covering all Transformer-based model families, targeting settings where only consumer grade computing resources are available.
+The work by He et al.
+[13] stands out among exiting descriptive studies, comprehensively addressing the capabilities, limitations, development and integration of language models in healthcare.
+The language models in scope are pre-trained and large language models.
+The development process is explained in detail, covering aspects such as training data, methodologies, and optimization strategies.
+Concerns related to the integration of LLMs into healthcare are also investigated, as fairness, accountability, transparency, and ethics.
+Zhou et al.
+[84] also provide a comprehensive overview of the development and deployment of LLMs in medicine, together with the challenges and opportunities these models face.
+Their study is both discursive and practical, being one of its highlights.
+The authors detail the principles of existing medical LLMs, comprising basic model structures, number of parameters, and data sources and scales used for model development.
+A comparison of the performance of different LLMs across various medical tasks, also against state-of-the-art lightweight models, is also included.
+Continuing with practical reviews, Soni et al.
+[85] assessed the cost-effectiveness of pre-training and fine-tuning in BERT, BioBERT, Clinical BERT, and XLNet for medical question answering tasks.
+Their results indicate that BERT-based models exhibit superior performance when fine-tuned with mixed datasets (i.e., general and clinical domain data), highlighting a gap in well-generalizable medical QA datasets.
+The results also suggest that initial fine-tuning on general domain datasets, such as SQuAD, before doing it on clinical datasets can enhance performance.
+Prompting techniques were not included in their evaluations.
+In a similar vein, Jahan et al.
+[45] studied the impact of data size for fine-tuning and that of prompts in zero-shot learning on model performance.
+Four large language models are evaluated on six benchmark biomedical text processing tasks across datasets.
+Zero-shot LLMs outperform state-of-the-art fine-tuned models, such as BioBERT, BioGPT, and BioBART, when fine-tuning data is scarce.
+As the amount of fine-tuning data increases, so does the performance of these state-of-the-art fine-tuned models, surpassing zero-shot LLMs.
+The study also highlights LLMs’ sensitivity to prompts, as variations in these led to significant differences in outcomes.
+No single LLM consistently excelled across all datasets and tasks.
+The authors advocate the training of biomedical LLMs on domain-specific corpora while recognizing LLMs’ potential for biomedical applications that lack large annotated data.
+Lehman et al.
+[86] further explored whether LLMs trained primarily on general web text are suitable for highly specialized, safety-critical domains such as medicine, or if domain-specific models are a better alternative.
+A total of language models, ranging from million to billion parameters, are evaluated on three clinical tasks.
+As part of the experiments, T5 models were trained from scratch using MIMIC-III and MIMIC-IV clinical notes to investigate the efficiency of clinical tokens.
+Their findings suggest that relatively small, specialized clinical models significantly outperform all in-context learning approaches, even when fine-tuned on limited annotated data.
+Neither the models’ ability to handle long texts nor decoder-only and instruction-tuned models are accounted for in their work.
+Lastly, Li et al.
+[87] focuses on pre-trained language models for long clinical text.
+A core limitation of Transformer-based models is their substantial memory consumption, leading to performance degradation in long clinical texts.
+To overcome this limitation, the authors pre-trained Longformer and BigBird, two long-sequence Transformers, on a large-scale clinical corpus, extending the maximum input length from to .
+These models consistently and significantly outperformed ClinicalBERT and other short sequence Transformers across ten tasks.
+Long-sequence Transformers enriched with clinical knowledge are thus capable of learning long-term dependencies in long clinical texts according to the results.
+No generative tasks and solely encoder-only models are considered in their evaluations.
 4 Methodology
-The experiments involve two target tasks: text classification and conditional text generation.
-Specifically, text classification is addressed through three different approaches: (i) contextual embedding similarity, (ii) natural language inference (NLI), and (iii) multiple-choice question-answering (MCQA).
-These are explained in Section 4.1.2.
-As detailed below, each task is evaluated using diverse datasets, models, and performance metrics.
-The evaluation process comprises 53 models, which exhibit a size spectrum ranging from 110 million to 13 billion parameters.
-These models belong to three knowledge domains: general, biomedical or scientific, and clinical.
-Their characterization can be found in Table 1.
-OpenLLaMA 13B, Flan-T5-XXL, T5-V1.1-11B, and T0++ are run using float16 precision.
-In order to conduct a comprehensive analysis, we recognize the importance of evaluating the chosen models across a broad spectrum of medical specialties, ultimately aiming at a general assessment of medical knowledge.
-To achieve this, we used the transcriptions dataset.
-Nevertheless, also recognizing the potential of artificial intelligence in particular medical fields, we narrowed our focus in addition to radiology, one of the most promising fields for AI integration.
-Consequently, we opt for the MIMIC-CXR dataset along with its labeled counterpart, MS-CXR.
-| Small (S) | Medium (M) | Large (L) | XL | XXL | ||||||||||||
-| ID | Model | Size | ID | Model | Size | ID | Model | Size | ID | Model | Size | ID | Model | Size | ||
-| Encoder-only | General | m00 | BERT 0 [14] | 110 M | m01 | BERT 0 [14] | 340 M | - | - | - | - | - | - | - | - | - |
-| Biomedical | m02 | BiomedBERT | 110 M | m04 | BiomedBERT-large | 340 M | - | - | - | - | - | - | - | - | - | |
-| (abstracts + full text) 0 [71] | (abstracts only) 0 [71] | |||||||||||||||
-| m03 | BiomedBERT | 110 M | m12 | RoBERTa-MNLI ‡1 [81] | 355 M | |||||||||||
-| (abstracts only) 0 [71] | ||||||||||||||||
-| m05 | SciBERT 0 [70] | 110 M | ||||||||||||||
-| m06 | SapBERT 0 [82] | 110 M | ||||||||||||||
-| m07 | BioLORD-STAMB2-v1 0 [83] | 110 M | ||||||||||||||
-| m08 | BioLORD-STAMB2-v1-STS2 0 [83] | 110 M | ||||||||||||||
-| m09 | BioLORD-PMB 0 [83] | 110 M | ||||||||||||||
-| m11 | NLI-DeBERTa ‡1 [84] | 100 M | ||||||||||||||
-| Clinical | m10 | Bio+Clinical BERT 0 [75] | 110 M | - | - | - | - | - | - | - | - | - | - | - | - | |
+A series of experiments on medical text classification and conditional text generation are carried out to understand better the behavior of language models under resource-constrained settings, i.e., settings with consumer-grade computing resources.
+In total, language models are evaluated, whose size ranges from million to billion parameters.
+The selection of these models spans the general, biomedical, and clinical knowledge domains and includes the three families of Transformer-based models.
+Moreover, only open-source, smaller than B parameters models are considered.
+Details on the selected models are found in Table 1 and Appendix B.
+All experiments are performed using a Quadro RTX 8000 GPU and CUDA version 12.2.
+To guarantee that the selected models align with consumer-grade computing resources, models with more than billion parameters (i.e., OpenLLaMA 13B, Flan-T5-XXL, T5-V1.1-11B, and T0++) are run with float precision.
+By halving the floating-point precision, these and billion parameter model versions are still viable in computational resource-constrained settings.
+The three families of Transformer-based models are considered for the text classification task via different approaches (described in Section 4.1.2), whereas solely decoder-only models are used for the conditional text generation task.
+Transcriptions, MIMIC-CXR, and MS-CXR have been chosen as evaluation datasets.
+Transcriptions covers a broad spectrum of medical specialties, allowing a general assessment of medical knowledge.
+MIMIC-CXR and its labeled version, MS-CXR, enable testing focused on radiology, one of the most promising fields for AI integration, narrowing the evaluation to specialized medical knowledge.
+Small (S)
+Medium (M)
+Large (L)
+XL
+XXL
+ID
+Model
+Size
+ID
+Model
+Size
+ID
+Model
+Size
+ID
+Model
+Size
+ID
+Model
+Size
+Encoder-only
+General
+m00
+BERT 0 [16]
+110 M
+m01
+BERT 0 [16]
+340 M
+m11
+NLI-DeBERTa ‡1 [88]
+100 M
+m12
+RoBERTa-MNLI ‡1 [89]
+355 M
+Biomedical
+m02
+BiomedBERT
+110 M
+m04
+BiomedBERT-large
+340 M
+(abstracts + full text) 0 [66]
+(abstracts only) 0 [66]
+m03
+BiomedBERT
+110 M
+(abstracts only) 0 [66]
+m05
+SciBERT 0 [65]
+110 M
+m06
+SapBERT 0 [90]
+110 M
+m07
+BioLORD-STAMB2-v1 0 [91]
+110 M
+m08
+BioLORD-STAMB2-v1-STS2 0 [91]
+110 M
+m09
+BioLORD-PMB 0 [91]
+110 M
+Clinical
+m10
+Bio+Clinical BERT 0 [70]
+110 M
 Encoder-
-decoder |
-General | m14 | T5-V1.1-Base 2 [85, 86] | 220 M | m13 | BART Large-MNLI 1 [87] | 407 M | m15 | T5-V1.1-Large 2 [85, 86] | 770 M | m16 | T5-V1.1-3B 2 [85, 86] | 3.0 B | m17 | T5-V1.1-11B 2 [85, 86] | 11.0 B |
-| m18 | Flan-T5-Base †2 [15] | 220 M | m19 | Flan-T5-Large †2 [15] | 770 M | m20 | Flan-T5-XL †2 [15] | 3.0 B | m21 | Flan-T5-XLL †2 [15] | 11.0 B | |||||
-| m22 | T0 3B †2 [88] | 3.0 B | m23 | T0++ †2 [88] | 11.0 B | |||||||||||
-| Biomedical | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | |
-| Clinical | m24 | ClinicalT5-base 2 [89] | 220 M | - | - | - | m25 | ClinicalT5-large 2 [89] | 700 M | - | - | - | - | - | - | |
-| Decoder-only | General | - | - | - | m26 | GPT-2 Medium3 [90] | 355 M | m27 | GPT-2 Large 3 [90] | 774 M | m28 | GPT-2 XL 3 [90] | 1.5 B | m29 | Palmyra Base 5B 23 [91] | 5.0 B |
-| m41 | OpenLLaMA 3B 3 [92] | 3.0 B | m30 | Camel 5B†2 [93] | 5.0 B | |||||||||||
-| m42 | OpenLLaMA 3Bv2 3 [92] | 3.0 B | m31 | GPT-J 6B 23 [94] | 6.0 B | |||||||||||
-| m32 | Instruct GPT-J †2 [95] | 6.0 B | ||||||||||||||
-| m33 | Falcon-7B 23 [96] | 7.0 B | ||||||||||||||
-| m34 | Falcon-7B-Instruct †2 [96] | 7.0 B | ||||||||||||||
-| m35 | MPT-7B 23 [97] | 7.0 B | ||||||||||||||
-| m36 | MPT-7B-Instruct †2 [97] | 7.0 B | ||||||||||||||
-| m37 | LLaMA-7B 23 [25] | 7.0 B | ||||||||||||||
-| m38 | LLaMA 2-7B 23 [26] | 7.0 B | ||||||||||||||
-| m39 | Alpaca 7B †2 [98] | 7.0 B | ||||||||||||||
-| m40 | LLaMA 2-CHAT-7B †2 [26] | 7.0 B | ||||||||||||||
-| m43 | OpenLLaMA 7B 3 [92] | 7.0 B | ||||||||||||||
-| m44 | OpenLLaMA 7Bv2 3 [92] | 7.0 B | ||||||||||||||
-| m45 | OpenLLaMA 13B 3 [92] | 13.0 B | ||||||||||||||
+decoder
+General
+m14
+T5-V1.1-Base 2 [92, 93]
+220 M
+m13
+BART Large-MNLI 1 [94]
+407 M
+m15
+T5-V1.1-Large 2 [92, 93]
+770 M
+m16
+T5-V1.1-3B 2 [92, 93]
+3.0 B
+m17
+T5-V1.1-11B 2 [92, 93]
+11.0 B
+m18
+Flan-T5-Base †2 [17]
+220 M
+m19
+Flan-T5-Large †2 [17]
+770 M
+m20
+Flan-T5-XL †2 [17]
+3.0 B
+m21
+Flan-T5-XLL †2 [17]
+11.0 B
+m22
+T0 3B †2 [95]
+3.0 B
+m23
+T0++ †2 [95]
+11.0 B
 Biomedical
-/ Scientific |
-- | - | - | m48 | BioGPT 3 [74] | 347 M | m47 | GPT-2-PubMed Large 3 [99] | 774 M | m50 | Galactica 1.3B 3 [24] | 1.3 B | m51 | Galactica 6.7B 3 [24] | 6.7 B | |
-| m46 | GPT-2-PubMed Medium 3 [99] | 355 M | m49 | BioGPT-Large 3 [74] | 1.5 B | |||||||||||
-| Clinical | - | - | - | - | - | - | - | - | - | - | - | - | m52 | MedAlpaca 7b †2 [58] | 7.0 B | |
+Clinical
+m24
+ClinicalT5-base 2 [96]
+220 M
+m25
+ClinicalT5-large 2 [96]
+700 M
+Decoder-only
+General
+m26
+GPT-2 Medium3 [19]
+355 M
+m27
+GPT-2 Large 3 [19]
+774 M
+m28
+GPT-2 XL 3 [19]
+1.5 B
+m29
+Palmyra Base 5B 23 [97]
+5.0 B
+m41
+OpenLLaMA 3B 3 [98]
+3.0 B
+m30
+Camel 5B†2 [99]
+5.0 B
+m42
+OpenLLaMA 3Bv2 3 [98]
+3.0 B
+m31
+GPT-J 6B 23 [100]
+6.0 B
+m32
+Instruct GPT-J †2 [101]
+6.0 B
+m33
+Falcon-7B 23 [102]
+7.0 B
+m34
+Falcon-7B-Instruct †2 [102]
+7.0 B
+m35
+MPT-7B 23 [103]
+7.0 B
+m36
+MPT-7B-Instruct †2 [103]
+7.0 B
+m37
+LLaMA-7B 23 [27]
+7.0 B
+m38
+LLaMA 2-7B 23 [28]
+7.0 B
+m39
+Alpaca 7B †2 [104]
+7.0 B
+m40
+LLaMA 2-CHAT-7B †2 [28]
+7.0 B
+m43
+OpenLLaMA 7B 3 [98]
+7.0 B
+m44
+OpenLLaMA 7Bv2 3 [98]
+7.0 B
+m45
+OpenLLaMA 13B 3 [98]
+13.0 B
+Biomedical
+/ Scientific
+m48
+BioGPT 3 [69]
+347 M
+m47
+GPT-2-PubMed Large 3 [105]
+774 M
+m50
+Galactica 1.3B 3 [26]
+1.3 B
+m51
+Galactica 6.7B 3 [26]
+6.7 B
+m46
+GPT-2-PubMed Medium 3 [105]
+355 M
+m49
+BioGPT-Large 3 [69]
+1.5 B
+Clinical
+m52
+MedAlpaca 7b †2 [71]
+7.0 B
 4.1 Text classification
-For the text classification task, we evaluate two datasets using three approaches and their corresponding models.
-It is essential to mention that fine-tuning is not performed; however, we use zero-shot prompt-engineering.
-In the following, we provide a concise overview of the datasets, followed by an explanation of the approaches employed.
-We refer the reader to Appendix A for more details on the preprocessing steps and the data characteristics of the datasets.
+Text classification is addressed using the Transcriptions and MS-CXR datasets and three different approaches: (i) contextual embedding similarity, (ii) natural language inference (NLI), and (iii) multiple-choice question answering (MCQA).
+The contextual embedding similarity approach is intended for encoder-only models, the NLI approach for encoder-only and encoder-decoder models pre-trained for NLI, and the MCQA approach for encoder-decoder and decoder-only models..
+Model tuning is implemented through zero-shot learning.
+To analyze the impact of prompting on text classification performance, different prompts are applied during inference.
+These prompts, grouped into two sets, are defined according to the classification approach.
+The first set of prompts is used for contextual embedding similarity and NLI.
+Since neither of these approaches requires a prompt to work, its non-use is also included in the analysis.
+The second set of prompts is used for MCQA, approach that needs a prompt to work.
+Prompts from the second set are defined based on those most commonly used in instruction-tuning models for multiple-choice question answering tasks.
+Let be a text sample and be a class, not necessarily corresponding to .
+A prompt from the first prompt set, , is defined as a function of a prompt template and a label.
+For example, “This is an example of ”.
+The set is only applied to the classes.
+Meanwhile, a prompt from the second prompt set, , combines a prompt template (consisting of the prompt structure and a question), a text sample, and the classes.
+For example, “You are a doctor and have the following information about a patient from a chest x-ray: .
+What is the diagnosis?
+.
+(”.
+In this example, the prompt template consists of the question “What is the diagnosis?” and the prompt structure, which is the rest of the text.
+Prompts are presented in detail in Appendix C.
 4.1.1 Datasets
-Transcriptions is a multi-label collection of electronic health records covering several medical specialties.
-The preprocessing procedure for this dataset involves removing null entries, modifying the report’s format, and selecting and renaming labels.
-After preprocessing, the resultant dataset consists of 2074 samples and 29 labels.
-Model performance is evaluated using the AUC score.
-In particular, since the EHRs are used in their entirety, certain samples contain token vectors that exceed the maximum input length allowed for multiple models.
-There are even samples whose token vectors exceed 2048 tokens, the maximum input length supported among the models.
-Consequently, different strategies have been adopted for each approach to effectively address this challenge.
-The adopted strategies, which require splitting the sentence into parts, will produce different results than processing the entire sample at once due to the attention mechanisms.
-These approaches are detailed in Section 4.1.2.
-MS-CXR is a multi-class dataset comprising sections of X-ray reports, each accompanied by annotations made by a radiologist [100, 101, 102].
-The number of classes is eight, well distributed and represented in 718 unique text samples.
-The preprocessing procedure is more straightforward than for the other datasets and includes removing samples with missing information and duplicates.
-In contrast to transcriptions, there is no evidence of any sample exceeding the maximum input length allowed for any of the models.
-Model performance is evaluated using a set of metrics in their macro-averaged version to ensure a comprehensive assessment. The metrics are accuracy, F1-score, precision, and recall.
+The datasets evaluated in text classification are Transcriptions and MS-CXR. Each of these datasets is introduced below. Their preprocessing and characterization details are given in Appendix A.
+Transcriptions is a multi-label collection of electronic health records (EHRs) covering many medical specialties.
+Preprocessing is applied to the data, removing null entries, organizing the EHR format, and selecting the final set of labels.
+After all, samples and classes are available.
+Performance is measured by the AUC score since the dataset is multi-label.
+Due to the length of some EHRs, certain token vectors exceed the maximum input length allowed by some models.
+To cope with this limit, the input sequence is processed using a non-overlapping sliding window method [87], as detailed in Section 4.1.2.
+MS-CXR is a multi-class dataset composed of X-ray report sections, each accompanied by annotations made by a radiologist [106, 107, 108].
+There are unique samples representing eight well-distributed classes.
+Preprocessing of this dataset consists of removing samples with missing information and duplicates.
+Contrary to Transcriptions, no sample exceeds the maximum allowed input length for any of the models.
+Performance is measured by accuracy, F1-score, precision, and recall in their macro-averaged version to ensure a comprehensive assessment.
 4.1.2 Approaches
-Contextual embedding similarity is grounded in the cosine similarity between the contextual embeddings of the text under consideration and its corresponding label.
-The models employed are exclusively encoder-only, comprising 11 models subjected to evaluation.
-The final contextual or sentence embedding is determined using three distinct pooling strategies: the CLS-token embedding, average pooling of the token-level embedding, and maximum pooling of the token-level embedding.
-In addition, prompt-engineering is implemented for the label.
-The specific details of the proposed prompts can be found in Appendix B.
-In cases where the input tokens exceed the maximum input size defined by the model, the samples are split into smaller parts, processing each individually.
-The embedding of the sample varies depending on the pooling strategy implemented, as described below.
-In particular, the models used in this approach have a maximum input token size of 512 tokens.
-CLS pooling: The contextual embeddings of the constituent parts of a sample are obtained by CLS pooling.
-The contextual embedding of the sample is calculated as the element-wise average among the contextual embeddings of the constituent parts.
-Maximum pooling: The contextual embedding of each sample is determined as the element-wise maximum of the output token-level embeddings of the constituent parts.
-Average pooling: The average operator considers the number of non-padding tokens of the parts into which a sample’s text is segmented.
-Thus, the contextual embedding of each sample is calculated as the element-wise weighted average among the token-level embeddings of its constitutive parts.
-Natural language inference for text classification involves representing the text to be classified as a premise and the labels as hypotheses.
-Subsequently, a multi-classification task with three classes -entailment, neutral, and contradiction- is performed.
-Conceptually, this approach can be thought of as having classifiers, where corresponds to the number of labels or classes.
-The models employed are encoder-only (cross-encoders) or encoder-decoder architectures.
-How the output is processed varies depending on whether the dataset is multi-label or multi-class.
-The entailment and contradiction outputs are transformed into binary probabilities in the multi-label scenario.
-This transformation effectively translates the problem into binary text classifiers, each classifier for one of the labels.
-In contrast, in the multi-class scenario, only the entailment logits are used to predict the class associated with a sample.
-When the length of the input tokens exceeds the maximum input size defined by the model, the scores of these samples are calculated by aggregating the scores obtained from their constituent parts.
-Aggregation measures that can be considered include maximum, average, or median.
-Furthermore, like the previous approach, the influence of prompts is considered during the model evaluation.
-The prompts used in this context are identical to those employed in the contextual embedding similarity approach, and their details can be found in Appendix B.
-Multiple-choice question-answering framework enables encoder-decoder and decoder-only architectures to perform text classification.
-A total of 27 models are assessed in this approach.
-Table 1, which provides details of the models used, shows the inclusion of pre-trained models and their instruction-tuned counterparts.
-For this approach to work, a prompt is required.
-Therefore, the impact of the prompt is measured in the analysis.
-The prompts are formulated based on those most commonly used in instruction-tuning models for multiple-choice question-answering tasks.
-However, in these prompts, the question format is not restrained.
-For this reason, this study also considers the impact of the question itself.
-Additional information on the prompts used in this approach can be found in Appendix B.
-It is essential to mention that the logit space has been constrained to align with the set of viable alternatives presented in a multiple-choice scenario.
-For this, the token identifiers associated with the feasible response options are determined and used to filter the logit space.
-4.2 Conditional text generation
-We evaluate one dataset for the conditional text generation task, using perplexity as the evaluation metric.
-Perplexity is a measure of uncertainty on the value of a sample from a discrete probability distribution.
-If we have a tokenized sequence , then
-where is the log-likelihood of the -th token conditioned on the preceding tokens according to our model.
-Below, a concise overview of the dataset employed is provided, followed by an explanation of the approach adopted.
-A more detailed characterization of MIMIC-CXR and its preprocessing procedure is presented in Appendix A.
-4.2.1 Datasets
-MIMIC-CXR is composed of X-ray reports [103, 104, 102].
-As part of the preprocessing procedure, the relevant sections are extracted using the code provided by Johnson et al.
+Text classification is performed through (i) contextual embedding similarity, (ii) natural language inference, and (iii) multiple-choice question answering.
+Contextual embedding similarity is grounded in the cosine similarity between the contextual embeddings of the sample text and the classes.
+The contextual or sentence embedding is determined by three distinct pooling strategies: CLS-token embedding, average token-level embedding pooling, and maximum token-level embedding pooling.
+For this approach, encoder-only models are employed, with a total of models evaluated.
+These models have a maximum input token size of tokens.
+Therefore, the samples’ token vectors that exceed this limit are processed with the non-overlapping sliding window method.
+The fragments are aggregated according to the pooling strategy, as follows.
+CLS pooling: The contextual embedding of each fragment from a sample is computed as its CLS-token output embedding.
+These embeddings are then aggregated using the element-wise average to obtain the contextual embedding representing the sample.
+Maximum pooling: The contextual embedding of each fragment from a sample is computed by applying element-wise maximum at token level over the output embeddings.
+These embeddings are then aggregated using again the element-wise maximum to obtain the contextual embedding representing the sample.
+Average pooling: The contextual embedding of each fragment from a sample is computed by applying the element-wise average at the token level over the output embeddings.
+These embeddings are then aggregated using the element-wise weighted average to obtain the contextual embedding representing the sample.
+Average’s weights indicate the number of non-padding tokens in each fragment.
+Natural language inference is the task of determining whether a hypothesis is true (entailment), false (contradiction), or indeterminate (neutral) given a premise.
+When applied for text classification, the premise represents a test sample, and the hypothesis represents the classes.
+For multi-class datasets, the predicted label is calculated from the entailment logits of each hypothesized class.
+For multi-label datasets, the entailment and contradiction logits are transformed into binary probabilities, which indicate whether or not a particular hypothesized class is predicted.
+This could be viewed as having binary text classifiers, where is the number of classes.
+This approach employs encoder-only (cross-encoder) and encoder-decoder models.
+These models have a lower maximum input token size than some of the test samples.
+Therefore, the token vectors of these particular samples are processed with the non-overlapping sliding window method.
+They are divided into fragments, whose scores are calculated individually, and then these scores are averaged to get the score of the whole sample.
+Multiple-choice question answering enables generative models, i.e., encoder-decoder and decoder-only models, to perform text classification.
+A total of models are assessed in this approach, including both pre-trained models and their instruction-tuned versions.
+As multiple-choice question answering is not intended for a extensive number of choices, the Transcriptions dataset is evaluated using a reduced version with eleven classes instead of the 29 available.
+These eleven classes consist of the ten most frequent labels plus an “Other” class.
+The number of samples evaluated is not affected.
+Additionally, the models’ logit space has been constrained to align with the response options of a multiple-choice scenario and, thereby, allow for automated evaluation.
+The token identifiers associated with the feasible response options are determined and used to filter the logit space.
+4.2 Conditional text generation task
+Conditional text generation is assessed with the MIMIC-CXR dataset, using perplexity as the performance evaluation metric.
+Perplexity (PPL) is a measure of uncertainty on the value of a sample from a discrete probability distribution.
+Let be a tokenized sequence, then
+where is the log-likelihood of the -th token conditioned on the preceding tokens .
+Decoder-only models are employed for evaluation, with a total of models considered.
+Among these models is Galactica, whose tokenizer lacks special tokens.
+As a consequence, two scenarios are analyzed: the inclusion and the non-inclusion of the start-of-sequence (BOS) token.
+The BOS token is a special token typically used by generative models to indicate the start of a text.
+In the first scenario, this token is included during tokenization, and perplexity is calculated from the first token in the texts.
+When the model’s tokenizer does not have the BOS token predefined, such as Falcon-7B, it is then defined as the tokenizer’s first special token.
+In the second scenario, the BOS token is excluded, and perplexity is calculated from the text’s second token.
+4.2.1 Dataset
+The dataset evaluated in conditional text generation is MIMIC-CXR, introduced below. Details on its preprocessing and characterization are in Appendix A.
+MIMIC-CXR is an X-ray reports dataset [109, 110, 108].
+Relevant sections of these reports are extracted using the code provided by Johnson et al.
 [111, 112].
-Following this extraction process and the subsequent removal of null and duplicated samples, the resulting dataset contains samples.
-Furthermore, none of the samples exceed the maximum input size allowed for the proposed models.
-4.2.2 Approaches
-For this task, 20 decoder-only models are evaluated, as detailed in Table 1.
-Among these models, we find Galactica, whose tokenizer lacks special tokens.
-As a result, the analysis is conducted in two parts.
-In the first part, the Beginning of Sequence (BOS) token is included during tokenization, with perplexity calculated from the first token of the texts.
-The BOS token is a special token usually employed by generative models to indicate the start of a text.
-In cases where a model’s tokenizer does not have the BOS token predefined, such as the Falcon-7B, it is defined as its first special token.
-The second part of the analysis proceeds without the BOS token, and perplexity is determined from the second token of the texts, and so on.
+Subsequently, null and duplicate samples are removed, with the resulting dataset having samples.
+None of these samples exceeds the maximum input size allowed for the proposed models.
 5 Results and Discussion
-The main findings are outlined below, each supported by the figures and tables provided. These findings can be further verified by the figures in Appendix C.
-5.1 Model size: more parameters do not generally translate into better performance
-The experiments conducted for text classification yield findings that challenge the assertion that larger models consistently deliver superior performance.
-An analysis of the impact of the logarithm of size on model performance indicates insufficient evidence to conclude that the Pearson correlation between size and performance is statistically significant in all the cases (refer to Fig.
-2).
-This observed discrepancy may be attributed to several factors, including the relatively small size of the models considered, with a number of parameters up to billions rather than tens or hundreds of billions.
-Specifically, in contextual embedding similarity, the trend of performance improvement with increasing size is almost nonexistent.
-As depicted in Fig.
-3, for instance, models such as SapBERT (m06) and BioLORD (m07-m09), which excel in this approach, outperform models even three times larger in both datasets.
-When solely the performance deltas associated with increasing the number of parameters for the same type of model are examined, they are both positive and negative.
-For the transcriptions dataset, as indicated by the lighter crosses in Fig.
-3, BERT (m01) marginally outperforms BERT (m00), similar to the case of BiomedBERT-large (abstracts only) (m04) versus BioMedBERT (abstracts only) (m03) where positive deltas are also obtained.
-However, for the MS-CXR dataset, BERT (m00) surpasses BERT (m01) in all metrics, and BioMedBERT (abstracts only) (m03) outperforms BiomedBERT-large (abstracts only) (m04) in precision.
-Now, by comparing BioMedBERT (abstracts only) (m03) and BiomedBERT (abstracts + full text) (m02), performance gains are obtained when more training data are used.
-In natural language inference, while the results align with the expectation that larger models tend to lead to better performance, more experiments are needed to draw a solid conclusion.
-In the MS-CXR dataset, the difference between the most effective NLI-DeBERTa model and the least effective RoBERTa-MNLI and BART Large-MNLI models ranges from across the metrics considered.
-These differences suggest that larger models among the models used always lead to performance improvement.
-However, reaching the same conclusion for the transcriptions dataset is not straightforward, given the results especially obtained for RoBERTa-MNLI (m12), which, as depicted in Fig.
-3, is closer to the performance of NLI-DeBERTa (m11) than to that of BART Large-MNLI (m13).
-When it comes to multiple-choice question-answering, the impact of the increase in size on performance is also unclear.
-While positive Pearson correlations are obtained, there is insufficient evidence to deem them statistically significant.
-When analyzing this impact within the T5 model family, as shown in Fig.
-3, the effect is minimal or inconsistent when considering their non-instruction-tuned versions (m14-m17, m24-m25).
-Conversely, when examining the instruction-tuned versions of T5 (m18-m23), a consistent positive effect of size on performance is observed for the transcriptions dataset.
-Although a positive effect is also present for the MS-CXR dataset, it is only for T0 models (m22-m23).
-Overall, the results of the text classification task suggest that increasing model size only sometimes leads to improved performance, whether comparing different models or evaluating the same models.
-Therefore, while model size may be a relevant factor in determining performance, training data and objectives may be more decisive.
-Expanding the sample size could be essential to validate this observation, considering a minimum of 30 or 35 models per approach.
-The previous insights can be verified in more detail in Figs.
-13, 14, 15, 16, 17 and 18 in Appendix C.
-In the context of the conditional text generation task, as shown in Fig.
-4, increasing the size of the model contributes to higher performances, regardless of whether the BOS token is included.
-Also, an improvement in the performance of the second versions (m42, m44) versus the first versions (m41, m43) of OpenLLaMA is observed.
-Considering that the difference between these OpenLLaMA versions is the dataset used for pre-training, the conditional text generation task results do not contradict the above conclusion for text classification.
-Furthermore, our findings do not conflict with existing scaling laws, usually associated with LLMs.
-First, because we have no control over training, i.e., we cannot guarantee that the model is not exposed to variations other than the number of parameters (size).
-Second, Chinchilla’s law does not consider inter-model comparisons.
-5.2 Model domain: more than a specialized domain; model architecture, training data, and training objective
+The main findings are outlined below.
+For comparability, AUC scores reported in this section correspond to evaluating the eleven class-reduced version of Transcriptions (see section 4.1.2).
+In addition, to ensure the robustness of the results, bootstrapping with iterations is applied to each experiment.
+Supplementary results are found in Appendix D.
+5.1 Text classification analysis
+| Dataset | Metric | CES | NLI | MCQA | ||||||
+| Model | Score | Prompt | Pooling | Model | Score | Prompt | Model | Score | ||
+| MS-CXR | Accuracy | BioLORD-STAMB2-v1-STS2 | x | Avg. | RoBERTa-MNLI | x | T0++ | |||
+| F1-score | BioLORD-STAMB2-v1-STS2 | Avg. | RoBERTa-MNLI | x | T0++ | |||||
+| Precision | BioLORD-PMB | CLS | RoBERTa-MNLI | x | Alpaca 7B | |||||
+| Recall | BioLORD-STAMB2-v1-STS2 | Avg. | RoBERTa-MNLI | x | T0++ | |||||
+| Transcriptions | AUC score | BioLORD-STAMB2-v1-STS2 | x | Avg. | BART Large-MNLI | x | Flan-T5-XXL | |||
+The highest F1 and AUC scores are achieved with the largest instruction-tuned T5 models, i.e., Flan-T5 (m19-m21) and T0 (m22-m23), as shown in Table 2 and Fig.
+2.
+Some of these scores are above in the F1-score and in the AUC score.
+These instruction-tuned T5 models in question stand among all models considered, ranking within the top 10 highest-performing models in both datasets.
+Nevertheless, the optimal choice of models may vary when considering precision as the target metric, where Alpaca (m39) and LLaMA 2-CHAT-7B (m40) demonstrate high competence.
+Conversely, the lowest F1 and AUC scores are paradoxically obtained with the base T5 models (m14-m17), as evidenced in Fig.
+2.
+These models, along with their clinically fine-tuned versions (m24-m25), rank within the top 10 lowest-performing models on both datasets.
+Moreover, the (1/1) and (6/8) of the models underperforming a random evaluator in Transcriptions and MS-CXR datasets, respectively, belong to base and clinically fine-tuned T5 models.
+Delving into each approach, BioLORD models (m07-m09) are consistently the best choice for the contextual embedding similarity approach.
+MS-CXR dataset is relatively more complex than the Transcriptions dataset for these models, as reflected by their ranking in performance: 11th versus 3rd place, respectively.
+BART Large-MNLI (m13) represents the best overall model for the NLI approach.
+For both datasets, BART Large-MNLI is included in the top 10 highest-performing models, while RoBERTa-MNLI (m12) only does so for MS-CXR dataset.
+For the multiple choice QA approach, instruction-tuned models stand out, which include instruction-tuned LLaMa models (m39-m40) to the aforementioned instruction-tuned T5 models.
+Notably, LLaMA 2-CHAT-7B (m40) is within the top 10 highest-performing models in both datasets.
+These highest performers per approach consistently show results indicative of clinical knowledge or clinical notions.
+The results of instruction-tuned T5 models support the feasibility of representing discriminative tasks as generative ones by framing them as instructions.
+These results also underline that generative tasks are not exclusive to decoder-only models, and text-to-text models may be a promising architecture to explore further.
+For example, versions of T5 tuned to instructions with 3B parameters (m20, m22) provide superior results to decoder-only models, almost three times larger, on both evaluated datasets.
+Model size – More parameters alone do not always translate into better results
+The experiments yield findings questioning the claim that larger models consistently deliver superior performance.
+The performance of the models as a function of the logarithm of their size is depicted in Fig.
+3.
+Testing for monotonic relationships via Spearman’s correlation is only reported for the multiple-choice QA approach due to the lack of diversity in sizes or number of samples.
+There is insufficient evidence to conclude that the Spearman’s correlation between size and performance is statistically significant in either dataset.
+The trend of performance improvement with increasing size is almost nonexistent in the contextual embedding similarity approach.
+As seen in Fig.
+2, for instance, models such as SapBERT (m06) and BioLORD models (m07-m09), which excel in this approach, outperform even three times larger models on both datasets.
+Within the same models, the deltas in performance associated with increasing the number of parameters are inconclusive.
+BERT (m01) marginally outperforms BERT (m00) on the Transcriptions dataset, whereas the opposite is observed in all metrics on the MS-CXR dataset.
+BiomedBERT-large (abstracts only) (m04) surpasses, albeit marginally, BioMedBERT (abstracts only) (m03) on both evaluated datasets, excluding in precision.
+Furthermore, performance gains are evidenced when more training data is used, as shown by comparing BioMedBERT (abstracts only) and BiomedBERT (abstracts + full text) (m02).
+Similarly, the effect of increasing the size on performance is not sufficiently clear or strong in the multiple-choice question answering approach.
+While positive Spearman’s correlations are obtained, there is insufficient evidence to deem them statistically significant.
+None of the p-values are , so the null hypothesis that the two variables have no ordinal correlation cannot be rejected.
+Within T5 models, the effect is minimal or inconsistent when considering their non-instruction-tuned versions (m14-m17, m24-m25).
+Within the instruction-tuned T5 models (m18-m23), a consistent positive effect of size on performance is observed for both FlanT5 and T0 models on Transcriptions and only for the latter on MS-CXR.
+On the other hand, the results in NLI align with the expectation that larger models lead to better performance.
+However, more models are needed to draw a (solid) conclusion.
+To have a notion about the lower bound in performance between NLI-DeBERTa (m11) and the largest models evaluated, the difference between the highest and the lowest values obtained among the evaluated prompts, respectively, is calculated.
+Considering all metrics, these differences range between on MS-CXR, having thus that the largest models always lead to performance improvement.
+Reaching the same conclusion on the Transcriptions dataset is not straightforward, given the results obtained for RoBERTa-MNLI (m12), as depicted in Fig.
+2.
+This model’s performance is closer to the performance of NLI-DeBERTa than to that of BART Large-MNLI (m13).
+Altogether, the results do not provide sufficient evidence that only increasing the model size, in number of parameters, leads to an improvement in performance, whether comparing different or the same models.
+Although model size may be a relevant factor in determining performance, it is hypothesized that training data and objectives are more decisive in small pre-trained language models.
+This hypothesis aligns with findings in [10] and [12].
+Expanding the sample size and diversity could be essential to validate these observations, considering a minimum of 30 or 35 models per approach.
+Model domain – More than a specialized domain; model architecture, training data, and training objective
 Current medical datasets remain relatively small compared to those of the general domain, covering only a tiny region of the medical knowledge space [84].
-Since some of the available domain-specific models are specialized by continuous pre-training on only one of these datasets, their generalization ability may be hampered, counter to the goal of creating well-generalizable domain-specific models and even achieving the ideal of artificial general intelligence (AGI).
-As discussed previously, the results suggest that architecture, training data, and training objectives are crucial in determining model generalization, possibly outweighing model size.
-For the classification task through contextual embedding similarity, the effectiveness of domain specialization in improving performance is not evident, as displayed in Fig.
-3.
-For the transcriptions dataset, refer to the lighter blue crosses in the figure.
-For example, Bio+Clinical BERT (m10) achieves lower results than expected, being around the middle of the performance ranking.
-Even in biomedical domain specialization, BERT (m00), a general domain model, outperforms some of the BiomedBERT models (m02-m04), challenging the superiority of domain-specific models.
-These observations are valid for both datasets.
-In particular, the latter observation also extends to BERT (m01).
-Two comparisons arise when employing multiple-choice question-answering: T5 models (m14-m15) versus their clinical specialized versions (m24-m25) and Alpaca (m39) versus MedAlpaca (m52).
-However, the evidence supporting the hypothesis that domain specialization significantly improves performance is limited, as displayed in Fig.
-3.
-Specifically, in the transcriptions dataset, differences in AUC scores between ClinicalT5 (m24-m25) and T5 (m14-m15) are and for the base and large sizes, respectively.
-This trend persists across all prompts.
-Similarly, the difference in AUC scores between MedAlpaca (m52) and Alpaca (m39) is , negatively affecting performance.
-In the MS-CXR dataset, the difference in F1-scores between ClinicalT5 (m24-m25) and T5 (m14-m15) is and for the base and large sizes, respectively.
-For MedAlpaca (m52) versus Alpaca (m39), the F1-score deteriorates even by .
-Thus, the evidence suggests that domain specialization provides either minimal or nonexistent benefits for this task.
-Similar patterns to those discussed previously are identified in the text generation task, as observed in Fig.
+Domain specialization of models using only one of these datasets in question may limit their generalization ability [45].
+The effectiveness of domain specialization in improving performance is not evident in the contextual embedding similarity approach, as displayed in Fig.
+2.
+The domain-specific models considered in this approach are Bio+Clinical BERT (m10), BiomedBERT models (m02-m04), and SciBERT (m05).
+Bio+Clinical BERT achieves lower scores than expected, positioning around the middle of the performance ranking for this approach.
+Similarly, some of the BiomedBERT models are outperformed by BERT (m00) and BERT (m01), their general domain counterparts.
+These findings, present in both datasets, challenge the superiority of domain-specific models over general domain ones in the task being evaluated via contextual embedding similarity.
+Although existing, evidence supporting the effectiveness of domain specialization is still limited and unclear in the multiple-choice question answering approach.
+The models to be compared are T5 models (m14-m15) versus their clinical specialized versions (m24-m25), and Alpaca (m39) versus MedAlpaca (m52).
+Differences between ClinicalT5 and T5 models are and in AUC scores and and in F1-scores.
+Similarly, differences between MedAlpaca and Alpaca are in AUC scores and in F1 scores.
+Due to these values, it can not be clearly stated that domain specialization positively impacts performance.
+Considering the insights discussed and the remarkable performance of BioLORD (m07-m09) models, SapBERT (m06), Flan-T5 (m18-m21) models, and T0 (m22-m23) models in their respective approaches, the training data, training objectives, and model architectures are possibly critical in determining model generalization.
+Continual pre-training for named entity recognition or medical entity linkage using contrastive learning on UMLS data is likely one of the factors for the success of SapBERT and BioLORD models.
+Likewise, employing instruction-tuned text-to-text models represents a compelling approach to achieving high performance in multiple-choice QA.
+Due to the impossibility of concluding on the NLI approach, expanding the analysis to incorporate domain-specialized NLI models in biomedical and clinical domains could be valuable.
+Prompting and instruction-tuning key to model performance
+One of the central points of the study is to analyze the influence of prompting on the models and text classification approaches under investigation.
+Prompting impact is quantified as the difference in performance resulting from the prompt usage, with positive values indicating an improvement, in contextual embedding similarity and NLI.
+In multiple-choice QA, this impact is calculated as the variation in performance, expressed in standard deviations, when different instructions are used.
+The resulting distributions are shown in Fig.
 4.
-Remarkably, the second version of LLaMA, LLaMA 2-7B (m38), stands out as the model with the highest predictive capacity for the dataset, with an average perplexity of .
-In contrast, BioGPT (m48) presents the greatest difficulty in comprehending the dataset, with an average perplexity of .
-These results are paradoxical, especially considering that BioGPT (m48) is domain-specialized while LLaMA 2-7B (m38) is not.
-A more detailed analysis by size reveals that
-1.
-for the medium size, BioGPT (m48) does not outperform any of the general domain models, whereas GPT-2-PubMed Medium (m46) does;
-2.
-for the large size, domain specialization proves beneficial; and
-3.
-for the XL and XXL sizes, neither Galactica (m50-m51) nor BioGPT-Large (m49) clearly outperforms general domain models.
-Consequently, the only specialized models that demonstrate an advantage are GPT-2-PubMed (m46-m47).
-Considering the insights discussed and the remarkable performance of SapBERT (m06) and BioLORD (m07-m09) models in the contextual embedding similarity approach, as well as Flan-T5 (m18-m21) and T0 (m22-m23) in the multiple-choice question-answering approach, the fundamental role played by training data, training objectives, and model architectures is underscored.
-Continual pre-training for named entity recognition or medical entity linkage using contrastive learning on UMLS data might contribute to the success of SapBERT and BioLORD models in the contextual embedding similarity task.
-On the other hand, employing instruction-tuned text-to-text models represents a compelling approach to achieving high performance in multiple-choice question-answering.
-Additionally, exploring not only the hypothesis but also the impact of domain instruction fine-tuning could provide further insights into model performance and effectiveness in specialized domains.
-Given that it is impossible to conclude the impact of the model domain on NLI, expanding the analysis to incorporate domain-specialized NLI models in biomedical and clinical domains could be valuable.
-5.3 Prompting and instruction-tuned models: key in determining model performance
-One of the central points of the study is to analyze the influence of prompts on the models and text classification approaches under investigation.
-As shown in Fig.
-5, the results underscore the fundamental role of the prompt and its wording in the performance of the model, demonstrating both positive and negative impacts.
-Consequently, we advocate using prompts and advanced prompt engineering techniques to guide the model toward better results.
-Nevertheless, this process should not be limited to a single prompt due to the observed and well-known phenomenon of prompt brittleness [11].
-As well, instruction fine-tuning proves advantageous for the models.
-Detailed results for each classification approach are presented below.
-In contextual embedding similarity, using a prompt does not always confer benefits.
-For the transcripts dataset, the mean impact on the AUC score is points, with differences ranging from to points.
-Overall, using any of the proposed prompts improves performance for of the model + pooling strategy combinations.
-As for the BioLORD-PMB, BiomedBERT family, BERT, and SciBERT models, none of the suggested prompts led to AUC score improvements.
-Inversely, for the MS-CXR dataset, the prompt impact on performance is generally more positive, albeit with more variability.
-The mean impact, along with its range, is (, ) for accuracy, (, ) for F1-score, (, ) for precision, and (, ) for recall.
-Overall, employing any of the proposed prompts represents benefits for to of the model + pooling strategy combinations, depending on the metric.
-As for the BioMedBERT (abstracts only) and BiomedBERT-large (abstracts only), their performance scores, except for precision, improve with any of the proposed prompts, while for the BioLORD family and Bio+Clinical BERT, their scores are always hindered.
-When examining the prompt impact on NLI, more consistent benefits are observed than for contextual embedding similarity.
-Using any of the proposed prompts yields improvements in the transcriptions dataset.
-The mean impact is for BART Large-MNLI, for NLI-DeBERTa, and for RoBERTa-MNLI, with larger models profiting the most from its usage.
-On its side, in the MS-CXR dataset, using a prompt only sometimes results in gains, particularly for NLI-DeBERTa.
-For this model, the impact on the F1-score is , while for BART Large-MNLI and RoBERTa-MNLI, it is and , respectively.
-For NLI-DeBERTa, positive prompt impacts are only observed for precision.
-For NLI, there are certain prompts with a high positive impact, whereas others do not, mostly independent of the model.
-Regarding multiple-choice question-answering, the importance of prompts is evident, as they can significantly influence model performance.
-For example, in the transcriptions dataset, the proportion of models performing better than random guessing (AUC scores greater than ) can improve from to with appropriate prompts.
-Similarly, in the MS-CXR dataset, this proportion can change from to when measured by the F1-score (scores greater than ).
-Thus, its importance is highlighted not only by the high performance obtained but also by its variability, as reflected in Fig.
-5.
-This is further supported by Figs.
-17 and 18 in Appendix C.
-When analyzing the impact of templates versus questions, both components of the prompt (see Appendix B), it is found that the wording of templates has more influence on performance than the wording of questions, and no single prompt (template + question) works universally well for all models.
-More specifically, in the transcriptions dataset, the performance ranking between the T5 models and between the Flan-T5 and T0 models remains consistent independent of the prompt.
-Moreover, Flan-T5-XXL (m21) and T0++ (m23) consistently outperform the other models regardless of the prompt.
-On average, the impact of the template, measured by the standard deviation of the AUC score per model, is , while the effect of the question within a template is .
-For the MS-CXR dataset, it becomes evident that some models, such as T5 and Falcon-7B, are unsuitable for this approach and dataset.
-The prompt plays an even more critical role than in the transcriptions dataset, affecting the ranking of models.
-The sensitivity is such that, with certain prompts, instruction-tuned models yield similar results to their base counterparts.
-On average, the impact of the template is , while the effect of the question within a template is .
-Interestingly, one of the proposed role-based prompts works remarkably well for the instruction-tuned versions of the T5 family despite not being the optimal choice overall.
-Concerning instruction-tuning, these models generally outperform their non instruction-tuned counterparts.
+Using a prompt does not always confer benefits in contextual embedding similarity, as reflected by Fig.
+4.
+On Transcriptions, the average impact on the AUC score is points, with values ranging from to .
+Using any of the proposed prompts improves performance for of the model + pooling strategy combinations.
+In contrast, none of these prompts led to AUC score improvements for BioLORD-PMB, BiomedBERT models, BERT, and SciBERT.
+On MS-CXR, the impact of the prompt on performance is more positive on average, albeit with more variability.
+The average impact on the F1-score is points, with values ranging from to .
+Similar values are reported on accuracy, precision, and recall.
+Employing any of the proposed prompts represents benefits for the to of the model + pooling strategy combinations, depending on the metric.
+The performance of BioMedBERT (abstracts only) and BiomedBERT-large (abstracts only) is enhanced with any of the prompts, whereas the performance of the BioLORD models and Bio+Clinical BERT is hindered.
+More consistent benefits are observed than in contextual embedding similarity when examining the prompt impact in the NLI approach.
+On Transcriptions, any of the proposed prompts yields performance improvements, profiting larger models the most from its usage.
+The average impact on the AUC score is for BART Large-MNLI, for NLI-DeBERTa, and for RoBERTa-MNLI.
+On MS-CXR, using a prompt only sometimes results in gains, particularly for NLI-DeBERTa.
+For this model, the average impact on the F1-score is ; while for BART Large-MNLI and RoBERTa-MNLI is and , respectively.
+Moreover, positive prompt impacts are only observed on precision for NLI-DeBERTa.
+In both datasets, there are certain prompts with a high positive impact, whereas others do not, mostly independent of the model.
+Similarly, prompt importance is also evident in the multiple-choice question answering approach, given its significant observed influence on model performance.
+The proportion of models performing better than a random evaluator (AUC score ) on Transcriptions increases from to with appropriate prompts.
+Similarly, the proportion of better than a random evaluator (F1-score ) on MS-CXR rises from to .
+Prompting importance is thus highlighted not only by the high performance achieved but also by the brittleness of the models.
+The latter is reflected by the variability in Fig.
+4, and further supported by Figs.
+17 and 18 in Appendix D.
+Between datasets, the highest sensitivity to the prompt is found when evaluating Transcriptions, such that, with certain prompts, the instruction-tuned models yield similar results to their base counterparts.
+Overall, no single prompt works universally well for all models.
+Regarding instruction-tuning, these models generally outperform their non instruction-tuned counterparts.
 The instruction-tuned T5 versions, whether T0 or Flan-T5, in any size considered, exhibit superior performance than their base counterparts.
-In the transcriptions dataset, Flan-T5 either performs better or is on par with T0.
-However, in the MS-CXR dataset, T0 outperforms Flan-T5 in the 11B parameter version only when comparing the highest performances due to the considerable variability of T0++.
-Also, instruction-tuning improves performance consistently for the LLaMA group, whereas this is not always the case for the other generative models: MPT and GPT-J are exceptions in the transcriptions dataset and Falcon in the MS-CXR dataset.
-Overall, instruction-tuning represents a gain, with mean improvements of in AUC-score and in F1 score between the highest performances per model and in AUC-score and in F1 score between the mean performances for the transcriptions dataset and the MS-CXR dataset, respectively.
-More details on the prompt impact can be found in Figs. 20 and 20 in Appendix C.
-5.4 Text classification: instruction-tuned T5: models to go for
-| CES | NLI | MCQA | ||||||||
-| Dataset | Metric | Model (poolingprompt) | Score | Model (prompt) | Score | Model (prompt) | Score | |||
-| Accuracy | BioLORD-STAMB2-v1-STS2 | (avg5) | 69.65 | RoBERTa-MNLI | (5) | 76.50 | T0++ | (t03+q04, t03+q05) | 81.68 | |
-| F1-score | BioLORD-STAMB2-v1-STS2 | (avg ) | 69.34 | RoBERTa-MNLI | (5) | 78.25 | T0++ | (t03+q05) | 83.90 | |
-| MS-CXR | Precision | BioLORD-PMB | (CLS ) | 83.34 | RoBERTa-MNLI | (5) | 80.74 | Alpaca 7B | (t02+q03) | 85.85 |
-| Recall | BioLORD-STAMB2-v1-STS2 | (avg ) | 72.58 | RoBERTa-MNLI | (5) | 82.28 | T0++ | (t03+q04) | 89.18 | |
-| Transcriptions | AUC score | BioLORD-STAMB2-v1-STS2 | (avg1) | 89.01 | BART Large-MNLI | (4) | 80.75 | Flan-T5-XXL | (t01+q09) | 92.38 |
-As shown in Table 2 and Fig.
-3, for both datasets, the most favorable results are obtained with larger instruction-tuned T5 models, namely Flan-T5 (m18-m21) and T0 (m22-m23).
-The latter highlights the possibility of representing discriminative tasks as generative ones by framing them as instructions.
-It is also emphasized that generative tasks are not exclusive to decoder-only models, and text-to-text models might be an architecture that needs to be explored more and potentially promising.
-For example, instruction-tuned versions of T5 with 3B parameters (m20, m22) provide higher results than decoder-only models, almost three times larger, applying to both datasets.
-When explicitly comparing precision and recall values, which can be critical depending on the application, the optimal choice of models may vary.
-Alpaca (m39) and LLaMA 2-CHAT-7B (m40) demonstrate proficiency in precision, while Flan-T5-XXL (m21) and T0++ (m23) excel in recall.
-Also, BioLORD (m07-m09) models, with 110 million parameters, exhibit precision and AUC score differences equal to or less than 3 points compared to models with a number of parameters between 7 billion and 11 billion.
-With respect to the datasets, the MS-CXR dataset poses more challenges for the models than the transcriptions dataset.
-This difference in difficulty could be attributed to factors such as (i) the format of the data, (ii) the fact that the reports are from different medical subfields, and (iii) the length of the reports.
-Nevertheless, the highest performers per approach consistently show results indicative of clinical knowledge or clinical notions.
-5.5 The groups of generative models: LLaMA and GPT
-Two procedures are carried out to gain insight into whether the models exhibit similar perplexity behavior and to identify potential clusters among them.
+Instruction-tuning also improves performance consistently for the LLaMA models, whereas this is not always the case for other generative models: MPT and GPT-J are exceptions on the Transcriptions dataset and Falcon on the MS-CXR dataset.
+Overall, this tuning technique represents a gain, with an average increase of points in the AUC score and points in the F1-score.
+Summarizing, the results endorse the crucial role of the prompt and its wording in the model’s performance, with both positive and negative effects presented.
+Consequently, we advocate using prompts and advanced prompting techniques to guide the model toward better results.
+This process should also not be limited to a single prompt due to the observed and well-known phenomenon of prompt brittleness [13].
+Regarding instruction tuning, this technique proves to be beneficial for the models.
+More details on the prompt impact can be found in Figs.
+20 and 19 in Appendix D.
+5.2 Conditional text generation analysis
+LLaMA models (m38-m39) stand out as the ones with the highest predictive capacity among the models evaluated.
+Particularly, LLaMA 2-7B (m38) is the highest performer, with a mean perplexity of when including the BOS token and when not.
+LLaMA models are also notable for the low standard deviation of their mean, with approximate values of and depending on the BOS token usage.
+These standard deviations indicate higher confidence in the estimated value of the mean.
+Conversely, BioGPT models (m48-m49) are the models with the most significant difficulty in comprehending the dataset.
+BioGPT (m48), the lowest performer, presents a mean perplexity of when including the BOS token and when not.
+The variability on the mean of these models is among the highest observed, with approximate standard deviations of and depending on the BOS token usage.
+These results are paradoxical considering that BioGPT is domain-specific while LLaMA 2-7B is not.
+Similarly to previous findings for text classification, domain specialization does not necessarily imply surpassing general domain models.
+For the medium-size domain-specific models, it is observed that BioGPT (m48) does not outperform any of the general domain models, while GPT-2-PubMed Medium (m46) does.
+For the large size models, domain specialization proves beneficial; whereas for the XL and XXL sizes, neither Galactica (m50-m51) nor BioGPT-Large (m49) clearly outperforms general domain models.
+Consequently, the only specialized models that prove advantageous are GPT-2-PubMed (m46-m47).
+On the other hand, increasing the model size contributes to improved performances, regardless of whether or not the BOS token is included.
+A slight performance improvement is also observed for the second versions (m42, m44) versus the first versions (m41, m43) of OpenLLaMA.
+This improvement is on average and points on the perplexity for the 3B and 7B parameter versions, respectively.
+Considering that the difference between these versions of OpenLLaMA is the dataset used for pre-training, the results obtained for conditional text generation do not contradict those for text classification.
+Furthermore, the standard deviations of perplexity reveal the presence of exceptionally challenging samples for the models, that is, outliers, which is visually depicted in Fig.
+21 in Appendix D.
+Moderate outliers, above quantile by times the IQR, represent between and of the data, with BioGPT models having the highest percentages.
+Extreme outliers, above quantile by three times the IQR, make up between and of the data, with most models exhibiting percentages around and .
+Groups of generative models – LLaMA and GPT-2
+Two procedures are carried out to determine whether the models exhibit similar perplexity behavior and identify potential clusters among them.
 The first procedure involves calculating the correlations between the models.
-Spearman’s and Pearson’s correlations are considered, assessing monotonic and linear relationships.
-The second procedure involves dimensionality reduction via UMAP, followed by hierarchical clustering, represented in Fig.
-6 using dendrograms.
-The correlation analysis reveals that, in general, all models are positively related.
-This positive relation indicates that most samples have a similar relative difficulty level for the models.
-However, the BioGPT models are the only exceptions, showing low Spearman correlations compared to the other models and positioning them considerably far from the rest of the evaluated models.
-The observed behavior in terms of monotonic relationships remains consistent regardless of the use of the BOS token.
-Regarding linear relationships, groups of models are clearly observed, including LLaMA models, OpenLLaMA models, and GPT-2 models.
-Although the previous groups are somewhat expected, some unexpected associations are also evident, such as between Falcon-7B and MPT-7B and between Palmyra Base 5B and GPT-J 6B.
-These two groups have notably weak intergroup correlations.
-With respect to the LLaMA and OpenLLaMA models, their linear relationship weakens when the BOS token is used, indicating more pronounced performance disparities.
-Possibly, training data plays an important role in this context, as it is essentially their main difference [92].
-Further insights, provided by the correlations and supported by the dendrograms in Fig.
-6, reveal the existence of two main groups of models: the GPT-2 models and the LLaMA models.
-The analysis of the standard deviations reveals the presence of exceptionally challenging samples for the models, that is, outliers.
-This phenomenon is visually depicted in Fig.
-4, which illustrates the distribution of perplexity between different models.
-Overall, the perplexity distributions have considerable variance, with the LLaMA models demonstrating the slightest variation.
-More detailed analyses are required to estimate the importance of these outliers and possible patterns among them.
+Spearman’s and Pearson correlations are considered, assessing monotonic and linear relationships, respectively.
+The second procedure consists of dimensionality reduction via UMAP, followed by hierarchical clustering, represented by dendrograms in Fig.
+6.
+Both procedures reveal the existence of two main groups of models: the GPT-2 and the LLaMA models.
+In general, all models are positively correlated, indicating that most samples have a similar relative difficulty for these models.
+BioGPT models (m48-m49) are the only exception to this.
+Further looking at the Pearson correlations, clustering patterns are present, where groups such as the LLaMA, the OpenLLaMA, and the GPT-2 models are identified.
+Although these previous clusters are somewhat expected, some unexpected associations are also evident, such as between Falcon-7B and MPT-7B and between Palmyra Base 5B and GPT-J 6B.
+Moreover, linear relationships between the LLaMA and OpenLLaMA models weaken, interestingly, when the BOS token is used, indicating more pronounced performance disparities.Possibly, training data plays a role, as it is essentially their main difference [98].
 6 Conclusion
-This study comprehensively explored a range of language models with varying sizes and domain fields.
-These models, encompassing 52 in total, were tested in two fundamental natural language processing tasks: text classification and text generation.
-The size range of these models goes from 110 million to 13 billion parameters, which is relatively small compared to recent language models.
-Our findings, based on these specific models, have significant implications, particularly for researchers and organizations operating under resource-constrained contexts.
-In text classification, our research has explored the effectiveness of three distinct approaches: context embedding similarity, natural language inference, and multiple-choice question-answering.
+This study comprehensively explores small pre-trained language models with varying sizes, architectural families, and domains.
+These models, being considered, are tested for two fundamental medical natural language processing tasks: text classification and conditional text generation.
+The size of the models ranges from million to billion parameters, which is relatively small compared to recent language models but suitable for consumer-grade computing resources.
+Our findings have significant implications, particularly for researchers and organizations operating under computational resource-constrained settings.
+For the text classification task, three distinct approaches are explored: context embedding similarity, natural language inference, and multiple-choice question answering.
 BioLORD and SapBERT models have demonstrated remarkable performance in text classification via contextual embedding similarity.
-Additionally, the instruction-tuned versions of T5, Flan-T5 and T0, followed by the instruction-tuned versions of LLaMA, have exhibited good results in the multiple-choice question-answering approach.
-However, for the NLI approach, we have identified the need for extended exploration, particularly in the context of specialized domain models.
-This implies that further research is needed to fully understand the potential of NLI in text classification, especially in specialized domains.
+Similarly, the instruction-tuned versions of T5, Flan-T5 and T0, followed by the instruction-tuned versions of LLaMA, have exhibited outstanding results in the multiple-choice question answering approach.
+Flan-T5 and T0 are remarkably good in both general medical and radiology-specific knowledge assessments.
+To fully understand NLI models’ potential, further exploration of this approach is needed, particularly in specialized domains.
 A common thread running through our findings is the significance of the prompt in improving text classification performance across different datasets and approaches.
-The significance of prompts extends beyond mere performance gains; they present a viable alternative to the resource-intensive processes of training and fine-tuning language models, which are often associated with substantial financial and environmental costs, especially for large-scale models.
-Moreover, effective prompt engineering is essential to mitigate prompt brittleness, ensuring more robust and reliable outcomes.
-Since, in general, medical datasets remain relatively small and cover only a small region of the medical knowledge space [79], domain-specific models, specialized by continuous pre-training on only one of these datasets, might see their generalization ability hindered.
-This practice could be one of the explanations for the results obtained.
-The results also suggest that the architecture, training data, and training objectives are crucial in determining the model’s generalization abilities, possibly outweighing model size.
-Regarding the text generation task, two groups of models are identified based on the perplexities obtained on MIMIC-CXR.
-These groups consist of the GPT-2 models and the LLaMA models.
-Notably, the LLaMA models stand out due to their low perplexities with minimal variation.
+This significance extends beyond performance gains; they present a viable alternative to the resource-intensive processes of training and fine-tuning language models, which are often associated with substantial financial and environmental costs.
+Effective prompt engineering is also essential to mitigate prompt brittleness, ensuring more robust and reliable outcomes.
+As prompt brittleness is evidenced during the study, and given its importance, further exploration in this line of research is recommended.
+Medical datasets often remain relatively small and cover only a small region of the medical knowledge space [84], so domain-specific models specialized using these datasets might see their generalization ability hindered.
+This practice could explain, to some extent, the results obtained.
+The results also suggest that the architecture, training data, and training objectives are crucial in determining the model’s generalization abilities, possibly outweighing the relevance of model size as a single variable.
+For the conditional text generation task, LLaMA models stand out due to their low perplexities with minimal variation.
+Two groups of models are also identified based on the perplexities obtained in MIMIC-CXR: a group consisting of GPT-2 models and another of LLaMA models.
 Further research is needed to identify and understand the outliers within these results, as they could hold important insights.
 In conclusion, this research highlights the critical role of prompts in language model inference and reaffirms the effectiveness of instruction-tuned generative models in addressing downstream tasks.
-It also highlights the significance of model architecture, training data, and training objectives, potentially even more so than model size, in its generalization capacity.
-We advocate for further investigations into topics such as model calibration, i.e., how certain the model is about output, prompt tuning, and performance concerning issues like hallucinations and biases, among others.
-Such studies can lead to more effective and ethical applications of language models, particularly in sensitive fields like healthcare.
+It also underscores the relevance of model architecture, training data, and training objectives, potentially even more so than model size alone, in its generalization capacity.
+We advocate for further investigations into topics such as model calibration, i.e., how certain the model is about output, prompt engineering and tuning, and performance concerning issues like hallucinations and biases, among others.
+Such studies can lead to more effective and ethical applications of language models in healthcare.
+Extensions to include quantized models and more medical NLP tasks will be considered in further research.
+Quantification is an interesting and promising approach to making LLMs viable in consumer-grade computing resources.
 Appendix A Data
@@
 However, this dataset also includes longer texts, some reaching up to 307 words.
-Appendix B Prompts
+Appendix B Models
+| ID | Model | Type | Domain | Model size (no. parameters) | Input max. size (no. tokens) | Total inference time (seconds) | |||
+| m00 | BERT [16] | Encode-only | General | M | – | ||||
+| m01 | BERT [16] | Encode-only | General | M | – | ||||
+| m02 |
+Encode-only | Biomedical | M | – | ||||
+| m03 |
+Encode-only | Biomedical | M | – | ||||
+| m04 |
+Encode-only | Biomedical | M | – | ||||
+| m05 | SciBERT [65] | Encode-only | Biomedical | M | |||||
+| m06 | SapBERT [90] | Encode-only | Biomedical | M | – | ||||
+| m07 | BioLORD-STAMB2-v1 [91] | Encode-only | Biomedical | M | – | ||||
+| m08 | BioLORD-STAMB2-v1-STS2 [91] | Encode-only | Biomedical | M | – | ||||
+| m09 | BioLORD-PMB [91] | Encode-only | Biomedical | M | – | ||||
+| m10 | Bio+Clinical BERT [70] | Encode-only | Clinical | M | – | ||||
+| m11 | NLI-DeBERTa [88] |
+General | M | – | ||||
+| m12 | RoBERTa-MNLI [89] |
+General | M | – | ||||
+| m13 | BART Large-MNLI [94] | Encoder-decoder | General | M | – | ||||
+| m14 | T5-V1.1-Base [92, 93] | Encoder-decoder | General | M | – | ||||
+| m15 | T5-V1.1-Large [92, 93] | Encoder-decoder | General | M | – | ||||
+| m16 | T5-V1.1-3B [92, 93] | Encoder-decoder | General | B | – | ||||
+| m17 | T5-V1.1-11B [92, 93] | Encoder-decoder | General | B | – | ||||
+| m18 | Flan-T5-Base [17] |
+General | M | – | ||||
+| m19 | Flan-T5-Large [17] |
+General | M | – | ||||
+| m20 | Flan-T5-XL [17] |
+General | B | – | ||||
+| m21 | Flan-T5-XLL [17] |
+General | B | – | ||||
+| m22 | T0 3B [95] |
+General | B | – | ||||
+| m23 | T0++ [95] |
+General | B | – | ||||
+| m24 | ClinicalT5-base [96] | Encoder-decoder | Clinical | M | – | ||||
+| m25 | ClinicalT5-large [96] | Encoder-decoder | Clinical | M | – | ||||
+| m26 | GPT-2 Medium [19] | Decoder-only | General | M | – | ||||
+| m27 | GPT-2 Large [19] | Decoder-only | General | M | – | ||||
+| m28 | GPT-2 XL [19] | Decoder-only | General | B | – | ||||
+| m29 | Palmyra Base 5B [97] | Decoder-only | General | B | |||||
+| m30 | Camel 5B [99] |
+General | B | – | ||||
+| m31 | GPT-J 6B [100] | Decoder-only | General | 6.0 B | 2 048 | ||||
+| m32 | Instruct GPT-J [101] |
+General | B | – | ||||
+| m33 | Falcon-7B [102] | Decoder-only | General | B | |||||
+| m34 | Falcon-7B-Instruct [102] |
+General | B | – | ||||
+| m35 | MPT-7B [103] | Decoder-only | General | B | |||||
+| m36 | MPT-7B-Instruct [103] |
+General | B | – | ||||
+| m37 | LLaMA-7B [27] | Decoder-only | General | B | |||||
+| m38 | LLaMA 2-7B [28] | Decoder-only | General | B | |||||
+| m39 | Alpaca 7B [104] |
+General | B | – | ||||
+| m40 | LLaMA 2-CHAT-7B [28] |
+General | B | – | ||||
+| m41 | OpenLLaMA 3B [98] | Decoder-only | General | 3.0 B | 2 048 | – | |||
+| m42 | OpenLLaMA 3Bv2 [98] | Decoder-only | General | 3.0 B | 2 048 | – | |||
+| m43 | OpenLLaMA 7B [98] | Decoder-only | General | 7.0 B | 2 048 | – | |||
+| m44 | OpenLLaMA 7Bv2 [98] | Decoder-only | General | 7.0 B | 2 048 | – | |||
+| m45 | OpenLLaMA 13B [98] | Decoder-only | General | 13.0 B | 2 048 | – | |||
+| m46 | GPT-2-PubMed Medium [105] | Decoder-only | Biomedical | 355 M | 1 024 | – | |||
+| m47 | GPT-2-PubMed Large [105] | Decoder-only | Biomedical | 774 M | 1 024 | – | |||
+| m48 | BioGPT [69] | Decoder-only | Biomedical | 347 M | 1 024 | – | |||
+| m49 | BioGPT-Large [69] | Decoder-only | Biomedical | 1.5 B | 1 024 | – | |||
+| m50 | Galactica 1.3B [26] | Decoder-only | Biomedical | 1.3 B | 2 048 | – | |||
+| m51 | Galactica 6.7B [26] | Decoder-only | Biomedical | 6.7 B | 2 048 | – | |||
+| m52 | MedAlpaca 7b [71] | Decoder-only | Clinical | B | – | ||||
+Appendix C Prompts
 The prompts used for the text classification task via contextual embedding similarity, natural language inference (NLI), and multiple-choice question answering (QA) are presented.
-B.1 Prompts for text classification via contextual embedding similarity and NLI
+C.1 Prompts for text classification via contextual embedding similarity and NLI
 The prompts proposed for text classification using contextual embedding similarity and Natural Language Inference (NLI) are exclusively applied to the label (in the case of NLI, to the hypothesis).
 Table 4 lists the prompts used.
-Template ID 0 is the default to generate the hypothesis in the zero-shot text classification using the NLI setting, as documented in HuggingFace.
-| ID | Label prompt | Dataset |
+Prompt template ID 0 is the default to generate the hypothesis in the zero-shot text classification using the NLI setting, as documented in HuggingFace.
+| ID | Prompt template | Dataset |
 | 0 | This example is {label}. | Transcriptions, MS-CXR |
@@
 | 7 | These findings are consistent with {label}. | MS-CXR |
-B.2 Prompts for text classification via multiple-choice QA
-The proposed prompts for text classification via multiple-choice question answering are based on the default templates specific to various of the considered instruction-tuned models.
+C.2 Prompts for text classification via multiple-choice QA
+The proposed prompts for text classification via multiple-choice question answering are based on the default prompt templates specific to various of the considered instruction-tuned models.
 These templates are systematically assessed using a set of questions, enabling us to quantify the influence of the question wording.
@@
 | q09 | Which one is the topic of the report? | Transcriptions |
-| ID | Prompt | Requirements | Dataset |
+| ID | Prompt structure | Requirements | Dataset |
 | t01 | Below is an instruction that describes a task, paired with an input that provides further context.
@@
 ( | report, options | MS-CXR |
-Appendix C Supplementary results
-This appendix presents supplementary figures and tables to support the results presented. They are displayed first by dataset and then by task or approach.
-C.1 Classification task
-C.1.1 Contextual embedding similarity
+Appendix D Supplementary results
+This appendix presents supplementary figures and tables to support the results presented.
+They are displayed first by dataset and then by task or approach.
+These results do not are not obtained by bootstrapping, but singular values for the complete inference dataset.
+D.1 Text classification task
+D.1.1 Contextual embedding similarity
 Results are depicted in Figs. 2, 14 and 13. The mapping between the models and their ID is
@@
 | m05: SciBERT |
-C.1.2 Natural language inference
+D.1.2 Natural language inference
 Results are depicted in Figs. 16 and 15. The mapping between the models and their ID is
 | m11: NLI-DeBERTa | m12: RoBERTa-MNLI | m13: BART Large-MNLI |
-C.1.3 Multiple choice question answering
+D.1.3 Multiple choice question answering
 | m14: T5-V1.1-Base | m21: Flan-T5-XXL | m31: GPT-J 6B | m38: LLaMA 2-7B |
@@
 | m20: Flan-T5-XL | m30: Camel 5B | m37: LLaMA-7B |
+D.2 Conditional text generation task
+Results are depicted in Fig. 21. The mapping between the models and their ID is
+| m26: GPT-2 Medium | m35: MPT-7B | m44: OpenLLaMA 7Bv2 | m50: Galactica 1.3B |
+| m27: GPT-2 Large | m37: LLaMA-7B | m45: OpenLLaMA 13B | m51: Galactica 6.7B |
+| m28: GPT-2 XL | m38: LLaMA 2-7B | m46: GPT-2-PubMed Medium | |
+| m29: Palmyra Base 5B | m42: OpenLLaMA 3B | m47: GPT-2-PubMed Large | |
+| m31: GPT-J 6B | m42: OpenLLaMA 3Bv2 | m48: BioGPT-Large | |
+| m33: Falcon-7B | m43: OpenLLaMA 7B | m49: BioGPT-Large |